### v1

# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

Abstract.
Extracting key information from documents represents a large portion of business workloads and therefore offers a high potential for efficiency improvements and process automation. With recent advances in deep learning, a plethora of deep learning-based approaches for Key Information Extraction have been proposed under the umbrella term Document Understanding that enable the processing of complex business documents. The goal of this systematic literature review is an in-depth analysis of existing approaches in this domain and the identification of opportunities for further research. To this end, 96 approaches published between 2017 and 2023 are analyzed in this study.
Key Information Extraction, Document Understanding, Business Documents, Deep Learning, Systematic Literature Review
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURccs: Applied computing Document analysisccs: Applied computing Business process management systemsccs: Computing methodologies Information extractionccs: Computing methodologies Neural networks
1. Introduction
The general idea of a paper-free – or at least paperless – office already came up five decades ago (Giuliano, 1975). However, to this day, physical paper documents still play an important role in business operations, as they are a key means of communication related to transactions both within and between organizations (Skalický et al., 2022). The processing of such documents is an essential yet time-consuming task that offers a high potential for automation due to the high workload involved as well as the critical nature of information transfer between different information systems (Cui et al., 2021; Voerman et al., 2021). At the same time, it can be observed that the ongoing digital transformation of business operations is leading to an increase in the digital processing of documents. This trend reinforces the need – but also the potential – for automated document processing, as more and more documents are available in digital form (Saifullah et al., 2023a).
Research on document processing is not new and has been conducted for several decades (Klein et al., 2004). In fact, the term ”document analysis” can be traced back to the 1980s (Wong et al., 1982). In recent years, however, there has been an upsurge in research related to document processing based on visually-rich documents (VRDs) and business documents, made possible by major advances in Deep Learning (DL). One of the most studied document processing task in this regard is Key Information Extraction (KIE) (Martínez-Rojas et al., 2023), which is concerned with extracting specific named entities from documents in a structured form (Skalický et al., 2022).
Complex business documents pose a significant challenge to KIE systems because they cannot be understood and processed as linear text sequences, as has been the case in most traditional KIE applications (Borchmann et al., 2021). These documents typically contain implicit and explicit cues such as different font formatting to distinguish headers from other parts of the document or complex positional dependencies between certain text segments. In this regard, related research investigates different methods to integrate such cues into the model architectures in order to achieve better extraction results. Business documents also have other special characteristics resulting from their connection to business processes. For example, different documents that are being processed as part of the same process run usually contain reoccurring information. These aspects and how they can improve corresponding KIE systems are worth investigating. In general, a process-oriented understanding is critical in order to adequately address the challenge of KIE in real-world contexts.
Although many DL-based KIE approaches for VRDs have been proposed, there is no comprehensive overview of most recent work in this area that focuses on the underlying DL concepts and technical characteristics while also adopting a business process perspective. The aim of this systematic literature review (SLR) is to fill this gap and provide a detailed overview of this research area and its state of the art. The contribution of this work is threefold:
  1. (1)
    A SLR based on 96 approaches to provide a concise overview of DL-based KIE methods for business documents.
  2. (2)
    Categorization and detailed comparison of corresponding methods based on various characteristics.
  3. (3)
    Dissemination of results, research gaps and derived potentials for future research.
The remainder of this manuscript is structured as follows: Chapter 2 provides background information on key concepts and nomenclature that are relevant to this SLR. In chapter 3, we discuss related work in terms of existing surveys and illustrate, how this study differs. The methodology used for this SLR, and more specifically how relevant literature was identified, is illustrated in chapter 4. Chapter 5 provides an overview of the results of the in-depth analysis. Chapter 6 serves as a discussion of the key findings and is dedicated to specific aspects of the identified literature. Based on the previous two chapters, we propose a research agenda and starting points for follow-up research in chapter 7. Chapter 8 concludes the manuscript with a summary.
2. Background
2.1. A business process perspective
Although business documents have a value of their own, they are best understood in relation to each other. Whenever they belong to the same process run, the information contained in these documents is also closely related to each other. For example, all documents related to the same run of a purchasing process typically contain the same order ID. Because they are embedded in specific processes that specify what information is relevant, business documents always contain distinct sets of predefined entities (Cristani et al., 2018). For example, invoices contain many different types of monetary values and unique identifiers such as invoice numbers that need to be identified.
Understanding temporal relationships in business processes is also important. Consider the following example of a simplified purchasing process. First, a customer places an order for a physical good. The company then processes the order, which results in the company sending an order confirmation and, at a later stage, the actual physical goods, a delivery note and an invoice. Figure SF1 of the electronic supplementary material shows an exemplary run of a corresponding process, including the interactions with internal and external entities such as customers and suppliers. Sequences where business documents are relevant are highlighted in red, indicating parts of the process where Document Understanding approaches could play a key role. As can be seen, especially interfaces related to external partners are expressed through document-based exchanges.
In addition to implicit knowledge about predefined entities, the consideration of associated business processes can also be useful in the sense that the information flow between individual process steps could be taken into account. For example, one could consider how the data extracted from a business document is further processed in subsequent process steps. From this, conclusions could be drawn as to how corresponding business documents should be processed. One could also analyze whether, in addition to the documents themselves, there is an additional data flow between the document processing steps that could facilitate the understanding of the documents.
2.2. Document Understanding
Document Understanding (DU), based on concepts from Natural Language Processing (NLP) and Computer Vision (CV), is an umbrella term covering a wide range of document processing tasks, including KIE, table understanding, document layout analysis (DLA), document classification and visual question answering (VQA) (Borchmann et al., 2021). More recently, novel tasks have been proposed such as ”Key Information Localization and Extraction”, which is an extension of KIE that emphasizes the need to localize key information, e.g., by identifying bounding boxes [^1] A bounding box indicates the coordinates of an element in the document image by its top left corner as well as its width and height. (Skalický et al., 2022). For a brief introduction to DU tasks besides KIE that are not covered in this review, we refer to (Cui et al., 2021). Other terms are also used for this research area, such as Document Analysis, Document AI or Document Intelligence (Motahari et al., 2021; Cui et al., 2021).
DU can also be grouped according to the complexity of the tasks, as suggested by (Du et al., 2022): Perception tasks deal with the recognition of descriptive document elements (e.g., text or entire tables), whereas induction tasks aim at the extraction of enriched information (e.g., document class or named entities) based on the perceived documents. Finally, reasoning represents the most complex subtask, combining perception and induction to enable document understanding beyond the explicitly contained information, mostly in the context of VQA. An example of a reasoning task is obtaining natural language explanations of figures and diagrams in documents (Skalický et al., 2022).
2.3. Key Information Extraction
KIE investigates methods for the automated extraction of named entities from documents into structured formats (Skalický et al., 2022) and can be further subdivided into individual research areas, namely Named Entity Recognition (NER), Named Entity Linking, Coreference Resolution, Relation Extraction (RE), Event Detection and Template Filling (Eisenstein, 2019; Oral et al., 2020), with NER and RE being the most prominent ones. The goal of NER is to detect entities in text and assign predefined labels to them, usually solved as a sequence labeling task (Balog, 2018). Figure 1 visualizes the outlined hierarchy of DU and subordinate research areas such as KIE. [^2] Note that this hierarchy is not a result of this study, but rather a reflection of the understanding in the literature.
[…]
A hierarchy of tasks and subtasks related to Document Understanding and Key Information Extraction
Figure 1. Overview of Document Understanding and Key Information Extraction
KIE approaches can be divided into three main subgroups based on how they represent the underlying documents, namely graph-based, grid-based and sequence-based methods. Graph-based systems convert document pages into graph structures that represent the layout and content of the pages. Such graphs typically allow for a flexible structure and can be constructed in a variety of ways. For example, each word or even character in the document image can be considered as a node in the graph. Different setups are also possible regarding the definition of the edges, for example creating a fully connected graph where every node is connected to every other node. Instead of constructing graphs, the grid-based approaches aim to create more organized grid structures with well-defined connections primarily characterized by rectilinear links – often based on the document image pixels. The grids are usually defined on different granularity levels, which affects which features the individual grid elements will be assigned to. For example, if the grid is defined on character level, all grid elements that are overlapped by the bounding box of a given character in the document image will have a value that is derived from that character (e.g., constant character index). Sequence-based techniques, on the other hand, convert documents into linear input sequences, ideally preserving and incorporating the document layout and other visual cues. These input sequences are then usually processed by Sequence Labeling methods, where each element is assigned to a particular class. For more details, we refer to (Sassioui et al., 2023). The following figure 2(c) illustrates the aforementioned paradigms and how they represent a sample document snippet.
(a) Graph-based (word-level)
(b) Grid-based (character-level)
(c) Sequence-based (word-level)
Figure 2. Main document representations of KIE methods (best viewed in color)
[…]
Different representations of a document snippet based on graph-based, grid-based and sequence based structures with different granularity levels
2.4. Input documents
In the context of DU research, input documents are often referred to as ”visually-rich documents”. Although there is no universally accepted definition, an aggregated understanding can be derived as follows: VRDs typically represent complex documents in which the simultaneous consideration of text, layout and visual information is of great importance in order to adequately capture their semantics (Cui et al., 2021; Xu et al., 2021; Katti et al., 2018; Gal et al., 2020). Examples of visual features are font properties such as bold text segments with an increased font size that represent a document title or specific keywords that indicate information to be extracted (Cui et al., 2021). Thus, all modalities are important for KIE and converting VRDs to linear text sequences would otherwise result in a significant loss of information. This is in contrast to ”simpler” documents such as news articles, where representations as linear text sequences are usually sufficient. This is also the reason why advances in DL enabled the processing of such complex input documents, as they allow for the integration and adequate processing of different input modalities.
The term ”business document” appears less frequently in related literature and there is also no universal definition of such documents (Cristani et al., 2018). In general, business documents contain process-relevant details related to the internal and external operations of an organization, as these documents represent a central means of communication (Cui et al., 2021; Skalický et al., 2022). Business documents, like VRDs, pose a significant challenge to automated DU systems for a variety of reasons, as discussed by (Motahari et al., 2021). For example, corresponding documents exist in a variety of different formats and are often only available in scanned form due to their paper-based distribution. The layout and overall content of business documents can vary from highly structured to highly unstructured. Furthermore, business documents can have relationships to other documents and/or may consist of multiple documents in a hierarchical fashion. To understand business documents, it is therefore necessary to observe such interrelationships and to understand the temporal relationships in processes.
As mentioned at the beginning, document processing is a central activity in business contexts. Therefore, it is promising to study KIE from a business process perspective and to investigate methods that directly address this challenge. The scope of this work regarding input documents is the intersection of VRDs and business documents. Therefore, we do not consider VRDs that are not typically embedded in business contexts, and at the same time, we do not consider business documents without characteristics of VRDs. An example of the latter is a company’s general terms and conditions. Although it can be considered as a business document, it typically appears as a simple text document without enriched visual elements. We use the terms VRDs and business documents interchangeably when addressing this intersection.
3. Existing reviews on Key Information Extraction
Many surveys are either domain-specific (e.g., healthcare (Wang et al., 2018)) or cover approaches that deal only with linear text inputs (Li et al., 2022a). In recent years, a few surveys have been published that cover DL-based DU methods while also dealing with VRDs. However, not all of them focus on the task of KIE (e.g., (Binmakhashen and Mahmoud, 2019) cover DLA).
In the following, we discuss the most closely related work. (Subramani et al., 2020) provide a brief KIE section that is limited to discussions on a few relevant aspects. The survey also does not include an in-depth analysis of the approaches. In (Baviskar et al., 2021), the authors provide a comprehensive overview of document processing with many different facets. However, it offers limited discussion about KIE. Much refers specifically to NER, where the overall workflow as well as different methods regarding pre-processing and feature extraction are being discussed. The authors discuss some model architectures. Since the chosen search period includes work between 2010 and 2020, a large number of analyzed papers also do not necessarily use DL methods and/or consider VRDs. (Cui et al., 2021) cover a wide range of DL-based DU topics. KIE, on the other hand, is discussed relatively briefly. The survey discusses aspects related to transformer-based document processing methods from a general point of view. However, details about the employed model architectures and other KIE paradigms are not covered in detail. The survey of (Antonio et al., 2022) focuses on preliminary DU tasks such as text detection and text recognition. KIE is also covered, although only a small number of approaches are included. The authors present some individual approaches in detail, however there is a lack of comprehensive overview. In (Martínez-Rojas et al., 2023), document processing is positioned in the context of Robotic Process Automation. Although an extensive literature survey is conducted, there is no in-depth technical discussion of KIE approaches for VRDs. An overview of key methods for KIE is provided by (Li et al., 2023a), albeit at a relatively high level. The authors however provide an extensive outlook on future work. Besides KIE, (Sassioui et al., 2023) also cover VQA and document classification. The authors propose a taxonomy for DU along different dimensions and focus on KIE benchmark datasets. Some challenges as well as future work are also discussed. (Abdallah et al., 2024) present a review of transformer-based methods for DU tasks. They highlight key paradigms of corresponding models and showcase a few approaches in detail. A major focus of the review is the description of benchmark datasets and related performance comparisons. In general, the survey chooses to discuss a few transformer-based approaches in detail, but does not present a broad overview of DL-based approaches with respect to KIE.
Overall, existing surveys examine KIE from a very broad perspective, cover only a few approaches in detail, or alternatively include a higher number of published work, but at the expense of a detailed discussion of the underlying methods. This work on the other hand provides an in-depth analysis, both quantitatively and qualitatively. One key distinguishing factor is also the adoption of a business process perspective. As discussed in section 2.1, the consideration of business processes as well as general domain knowledge is of high importance for KIE systems. To this end, only the review by (Martínez-Rojas et al., 2023) also adopts a practice-oriented perspective during the analysis. Table 1 summarizes the differentiation of the previously mentioned related work along certain distinguishing factors.
Table 1. Comparison against existing reviews
Distinguishing factor (Subramani et al., 2020) (Baviskar et al., 2021) (Cui et al., 2021) (Antonio et al., 2022) (Martínez-Rojas et al., 2023) (Li et al., 2023a) (Sassioui et al., 2023) (Abdallah et al., 2024) Ours
Search period n/a 2010-2020 n/a n/a 2017-2022 n/a n/a 2014-2023 2017-2023
Detailed search strategy ***** ***** **** *****
Focus on KIE task ** *** *** *** ** ***** *** *** *****
Business process perspective ** * **** *****
Analysis of DL concepts * *** ** ** * ** ** **** *****
Analysis of key characteristics ** *** **** *** *** **** **** *** *****
Analysis of KIE datasets * * *** * * **** *** ***** *****
Performance comparison ** ** * ** **** *****
Consideration of future work *** ** * ** *** *** * *****
4. Methodology
4.1. Research questions
The overall research question that is to be answered by this SLR is: ”What is the state of the art in Deep Learning based Key Information Extraction from business documents?”. To this end, eight further research questions have been defined:
  • RQ1: Which input modalities are considered and how are they integrated?
  • RQ2: Which DL architectures are being used?
  • RQ3: According to which criteria can existing approaches be categorized?
  • RQ4: Which input documents are considered?
  • RQ5: To what extent are practical applications and domain knowledge discussed?
  • RQ6: Which are the best performing approaches?
  • RQ7: Are there noticeable trends in the proposed approaches and architectures?
  • RQ8: What potential for improvement can be formulated for follow-up research?
The first five research questions cover an in-depth analysis of the identified approaches, with the aim of examining the proposed methods and how they differ from each other. This includes key aspects such as input modalities, model architectures and data bases. Based on this, research questions 6 to 8 adopt a more aggregated view and aim to identify the state of the art and derive recommendations for follow-up research.
4.2. Search procedure
The following six databases were used to obtain relevant literature: ACM Digital Library (ACM), ACL Anthology (ACL), AIS eLibrary (AIS), IEEE Xplore (IEEE), ScienceDirect (SD) and SpringerLink (SL). The search strings were carefully designed to include the commonly used terms for the research area as well as relevant keywords for the target domain (VRDs) and the application of DL architectures. If supported by the search engine, we also included terms that should not appear to filter out papers that are outside the scope of this work. The search period was limited to the range 2017 to 2023. Considering 2017 as the lower limit is a significant help to avoid false positive results, as KIE related literature typically did not use DL methods before 2017. The final search strings for each database, including the number of results at the time the query was run, can be found in table ST1 of the electronic supplementary material.
We defined the following inclusion and exclusion criteria as the basis for the literature screening. First, the title, abstract and conclusion were analyzed. If no violation of the criteria was identified on the basis of these parts, the full texts were analyzed subsequently.
  • IN1: The work is peer-reviewed and related to the research area of KIE.
  • IN2: The work employs DL concepts and outlines its architecture.
  • IN3: The work evaluates the effectiveness of the approach in a quantitative and/or qualitative manner.
  • EX1: The work is not written in English and/or was published outside of the defined search period.
  • EX2: The work does not propose a new approach for KIE, but rather applies existing approaches to different use cases, conducts a survey of existing approaches and/or only proposes a novel KIE dataset.
  • EX3: The work does not apply the approach to VRDs, but to other domains or only considers text-only inputs.
  • EX4: The work does not focus on KIE, but on other DU tasks and/or focuses only on image pre-processing tasks.
  • EX5: The work focuses only on extracting information from very specific document elements such as tables.
The overall search process is visualized in figure 3 according to the PRISMA defined by (Page et al., 2021). A total of 96 relevant approaches were identified for this SLR. A list of the identified approaches as well as their numbering, which is also used as a reference in this manuscript, can be found in table LABEL:mapping of the electronic supplementary material.
Figure 3. PRISMA flow diagram
[…]
Literature search procedure according to PRISMA flow diagram
5. Results
5.1. Overview
First, we provide a high-level overview of the analyzed work. To this end, table LABEL:overviewtab of the electronic supplementary material shows the results of the in-depth analysis along different properties, each of which considers different aspects, also with practical applications in mind. Figure 4(a) shows the overall distribution of the KIE paradigms. Almost half of the approaches belong to the group of sequence-based methods (n=46). Two further common categories are graph-based (n=24) and a combination of graph-based and sequence-based approaches (n=10). The grid-based representation is not as widely used as the graph-based representation, as only 9 approaches are purely based on this paradigm. Also, only 6 of the analyzed papers cannot be assigned to any of the groups, which underlines the dominance of the three paradigms for KIE systems. Interestingly, only the work by (Wang et al., 2021b) combines grid-based and sequence-based concepts into one KIE approach. Based on the distribution of the methods over time, as visualized in figure 4(b), one can see an increased interest in KIE research, especially since 2020. It is also noticeable that the dominance of sequence-based approaches first started in 2021, while graph-based methods were the most common group of approaches before that. The increased popularity of sequence-based methods in this period can probably be attributed to the influential work by (Xu et al., 2020), which was among the first to show state-of-the-art results using a transformer architecture and extensive pre-training procedures, which in turn led to a lot of follow-up research in this direction. One can also see a strong increase of graph-based methods in 2023 compared to the previous years as well as compared to the remaining categories. This category has therefore remained relatively popular over time. Overall, 2021 and 2023 saw the highest number of published papers. In 2022, however, there was a decrease in the number of corresponding KIE approaches (18 compared to 27 in the previous year). One reason for this observation may be the aforementioned increased interest in sequence-based methods, which typically require extensive pre-training. Therefore, many scholars spent 2022 on developing corresponding methods and on academic write-up, leading to a publication some time later in 2023.
(a) Overall distribution
(b) Distribution over time
Figure 4. Distribution of categories
[…]
A pie chart indicating the distribution of categories and a bar chart indicating the occurrences over time
The analysis of the integrated input modalities shows that the vast majority of approaches integrate at least textual (n=85) and layout-oriented (n=76) features. This is not surprising, since these elements contain the most relevant information for document processing. Layout modalities are usually derived from bounding boxes, i.e., the coordinates of words in document images. 56 of the analyzed approaches also integrate visual features obtained from the document images. As discussed in section 2.4, the integration of such visual cues into KIE systems is crucial for a proper understanding of complex VRDs. In this regard, one could even argue that, given this relevance, there are rather few approaches that integrate image modalities. Nevertheless, the results show that visual cues have became more and more important and have been increasingly integrated into the models over time, especially since 2021. In this regard, (Oral and Eryiğit, 2022) extensively investigate different methods to fuse visual modalities into KIE systems. Integrating multiple input modalities can play an important role for DU systems, as they incorporate cues from different aspects and therefore complement each other. For example, the image modality can provide relevant insights for documents with otherwise limited amounts of text and bounding boxes, while the layout modality is crucial for documents with very complex layouts and distinct structures (Yang et al., 2024). 36 of the approaches use hand-crafted input modalities. In most of these cases, they are integrated through custom input features such as boolean flags, whether a word in the document represents specific entities such as dates or monetary values (Palm et al., 2017; Gal et al., 2019; Lohani et al., 2019; Gal et al., 2020; Belhadj et al., 2021; Hamdi et al., 2021; Holeček, 2021; De Trogoff et al., 2022). Another feature type is information about the position of a particular word with respect to the overall reading order (Lee et al., 2021; Shi et al., 2023). In some papers, hand-crafted inputs are based on syntactic features such as the numbers of characters of a word (Sage et al., 2020; Hamri et al., 2023), details regarding the fonts (Wei et al., 2020; Hwang et al., 2021a) or encoded character representations (Krieger et al., 2021). (Zhang et al., 2023a) propose a feedback mechanism by fusing a label embedding based on decoder outputs back into the encoder for predicting adjacent fields. Therefore, the predicted label of an element also affects the prediction for adjacent elements. The use of hand-crafted features is also often associated with grid-based approaches. Such methods usually define a custom encoding function that assigns a specific value to each element of the grid. For example, (Palm et al., 2017; Dang and Thanh, 2020) assign a constant value to each pixel in a document image, such as an integer index of a character at each position – or the value 0 for blank parts of the document image. None of the approaches integrate meta data resulting from related business workflows as discussed in section 2.1, therefore lacking a practice-oriented perspective in this regard. Overall, hand-crafted features provide additional meta-level information beyond data that is explicitly contained in documents and can thus improve KIE systems. It must be said, however, that hand-crafted input features can potentially lead to a considerable amount of additional work in terms of labor-intensive data annotations (Saifullah et al., 2023a), which could also explain the less frequent use of this feature type. One could also argue that the inclusion of such hand-crafted features can potentially result in overfitting to the domain of the training documents – leading to poorer generalization capabilities across different domains. However, there is no work that analyzes this trade-off between integrating hand-crafted input features and the generalization capabilities across different domains.
In terms of the underlying data basis, the median number of employed documents is 2,702. This is a relatively low amount, especially since the data basis usually has to be split into training and evaluation partitions. As a result, many of the approaches are often only trained on a small document corpus and thus probably with little variety in terms of layouts. Since DL-based DU models usually require large amounts of data for training (Saifullah et al., 2023a), it is debatable whether the proposed approaches have reached their full potential. Note that these observations do not include the data basis that is being used for pre-training purposes, which will be discussed in the context of sequence-based approaches in section 5.2.1. Deviating from the average, (Palm et al., 2019; Klaiman and Lehne, 2021; Sarkhel and Nandi, 2021) employ more than 1 million documents for implementing their KIE system. On the other hand, in some cases, only 199 documents are used for developing the models, while still achieving promising extraction results (Li et al., 2021a, b; Du et al., 2022; Gemelli et al., 2023; Yu et al., 2023). The vast majority of used document types are receipts, forms and invoices. These three categories alone are used by around 70 % of the approaches. The main reason for this is the fact that the most frequently used benchmark datasets, namely CORD, FUNSD and SROIE [^3] These datasets will be discussed in more detail in section 5.3.1., are based on these document types. Invoices, on the other hand, are most often considered when authors do not use public benchmarks, but rather private in-house datasets.
Regarding the aspect of reproducibility, implemented code is available for 38 approaches. In most cases, authors share their code directly, however there are exceptions where implementations have been made available by external sources through re-implementation (e.g., the approach by (Powalski et al., 2021)). Model weights are shared even less frequently, which is however especially helpful for pre-trained sequence-based approaches. When either code or model weights are shared, they are mostly provided with a license that allows their application in commercial settings. This is beneficial for organizations that want to integrate corresponding KIE systems into their own business processes, which in turn helps to disseminate the research efforts.
10 approaches are independent of external OCR engines (Guo et al., 2019; Zhang et al., 2020; Wang et al., 2021a; Gu et al., 2022; Kim et al., 2022; Davis et al., 2023; Dhouib et al., 2023; Kuang et al., 2023; Yang et al., 2023a; Yu et al., 2023) and are therefore either responsible both for text reading and information extraction (end-to-end) or alternatively require no text reading stage at all and map input documents immediately to desired outputs (OCR-free) (Sassioui et al., 2023). The system by (Appalaraju et al., 2021) is only end-to-end during the pre-training phase. Also, (Klaiman and Lehne, 2021) propose a system that can be designed in an end-to-end manner, however the authors chose an external OCR engine in their work. These approaches first appeared in 2019, but only became more frequent in 2022 and 2023. Since this is a relatively new research direction, corresponding systems are relatively rare in the overall search period. The same is true for the 12 identified generative KIE methods (Klaiman and Lehne, 2021; Powalski et al., 2021; Wang et al., 2021b; Cao et al., 2022a, b; Kim et al., 2022; Cao et al., 2023; Davis et al., 2023; Deng et al., 2023a; Dhouib et al., 2023; He et al., 2023b; Tang et al., 2023), which first appeared in 2021, but now see an increased interest. The work by (He et al., 2023b) differs in the sense that the authors use an out-of-the-box Large Language Model (LLM) such as ChatGPT and prompt the model in specific ways including few-shot prompting to support KIE tasks. This in particular involves demonstrations for output formatting, but also making the models layout-aware in order to utilize positional relationships obtained by OCR. This category of KIE methods is able to output arbitrary texts using autoregressive decoding mechanisms. Key advantages are, that corresponding systems are not necessarily influenced by faulty OCR extractions as they can generate OCR-free representations of the target words (Cao et al., 2022b). They are also more flexible, as they can be conditioned of different DU tasks using distinct textual prompts (Kim et al., 2022; Davis et al., 2023). More traditional methods, which use a classification head for the downstream task such as KIE, would on the other hand need to be retrained with a different classification head in order to perform a different task (Davis et al., 2023).
Only 10 papers explicitly use domain knowledge (Palm et al., 2017; Gal et al., 2019; Palm et al., 2019; Majumder et al., 2020; Belhadj et al., 2021; Garncarek et al., 2021; Sarkhel and Nandi, 2021; Tata et al., 2021; Arroyo et al., 2022; Du et al., 2022). (Oral et al., 2020) also consider it, however only regarding their RE part of the approach and not the subsystem that is used for KIE. The integration of domain knowledge mostly consists of hand-crafted input features, as discussed before. Only two approaches have been evaluated in real-world industry settings, namely (Oral et al., 2020) and (Hwang et al., 2021a). More precisely, the authors show the impact of the developed KIE methods on real-world document processing tasks. 14 of the analyzed papers represent an evolution of an existing approach. A prominent example is the LayoutLM family. Based on the original LayoutLM, many subsequent refinements and improvements in various aspects have been proposed (Xu et al., 2020, 2021; Gu et al., 2022; Huang et al., 2022; Shi et al., 2023). Another example is Chargrid (Katti et al., 2018), which has been the basis for most other grid-based approaches (see section 5.2.3). One could argue that 14 out of 96 papers being an evolution is a relatively high proportion, and that as a result a lot of research is being done on iterative refinements and, to a lesser extent, disruptive innovations that could achieve significantly better KIE results. 8 of the analyzed approaches are implemented using weakly-annotated data. In most cases, this means that besides the document images, only the textual target values are required (Palm et al., 2019; Garncarek et al., 2021; Klaiman and Lehne, 2021; Wang et al., 2021b). Some other approaches only require annotations on segment-level or word-level (Ning et al., 2021; Cao et al., 2022b). Corresponding approaches therefore do not need to be trained on fully annotated documents, which usually involves a labeling of each word including its bounding box and entity.
The evolution of model sizes in terms of their number of parameters (in millions) is shown in figure 5. Note the necessary logarithmic scale of the y-axis due to the approach by (He et al., 2023b), which is based on GPT3 and therefore includes 175 billion parameters. The orange line indicates the trend over time, which shows a relatively small increase. Based on the analyzed approaches, the median number of trainable parameters is 157 million. Most of the approaches tend to have between 20 and 500 million parameters, which did not change significantly between 2017 and 2023. This is in stark contrast to the research field of LLMs in general, where models continue to grow exponentially with sizes often in the high billions.
Figure 5. Evolution of parameter counts
[…]
A chart that shows the evolution of parameter counts. The graph shows that there was no steep increase over time.
5.2. Categories
5.2.1. Sequence-based methods
Table LABEL:sequencetab of the electronic supplementary material presents the results regarding the analysis of sequence-based approaches. We also include hybrid KIE approaches that are based on multiple paradigms (e.g., a combination of graph-based and sequence-based methods) in this analysis and present the findings regarding the sequence-based subsystem. Most of the methods are defined on word-level (n=40), i.e., each word of a document image represents one element in the sequence. In some cases, a segment or sentence-level granularity is chosen (Gu et al., 2021; Li et al., 2021b, c; Ning et al., 2021; Tang et al., 2021; Wang et al., 2021a, 2022a; Kuang et al., 2023; Tu et al., 2023; Zhang et al., 2023a), which considers sentences as semantic entities and subsequently calculate aggregated visual and textual embeddings. Other possible granularities are on character-level (Guo et al., 2019; Yu et al., 2020), token-level (Wang et al., 2021a) or cell-level (Li et al., 2021a). Also, (Oral et al., 2020; Wang et al., 2021a, 2022a) use two different granularity types simultaneously, which allows the model to consider both granular and coarse features.
26 out of the 57 sequence-based approaches do not consider a pre-training of the models and subsequent fine-tuning steps for the KIE task. This is somewhat surprising, as it has been shown that obtaining a general document understanding through pre-training is a promising research direction. However, it should be noted that pre-training has been increasingly studied in the analyzed work since 2022. If pre-training is conducted, the average number of documents used is around 6 million. This highlights the relatively large amount of data that is required to pre-train corresponding models. The largest document corpus for pre-training is used by (Cao et al., 2022a), which consists of 43 million documents. Deviating from this is the approach of (Nguyen et al., 2021), which only uses 5,170 documents for pre-training procedures, while still achieving competitive benchmark results. Pre-training procedures are usually very resource-intensive. For example, (Xu et al., 2020) report that their largest model variant required 170 hours to finish one training epoch of a dataset consisting of 6 million documents. In around half of the cases, the IIT-CDIP dataset is used. This dataset has been proposed in (Lewis et al., 2006) and includes documents of the Legacy Tobacco Document Library [^4] https://www.industrydocuments.ucsf.edu/tobacco/ regarding lawsuits against the tobacco industry in the 1990s. Another common dataset is RVL-CDIP, which however represents a subset of IIT-CDIP and therefore includes no additional documents. Also, in six cases a private document collection has been used for pre-training procedures (Hwang et al., 2021a; Nguyen et al., 2021; Arroyo et al., 2022; Kim et al., 2022; Lee et al., 2022; Davis et al., 2023). Besides these common datasets, there exist other large document collections such as DocBank, which is however only considered by (Li et al., 2021c, 2022c; Tang et al., 2023; Yang et al., 2023a).
Table LABEL:sequencetab also lists the employed pre-training tasks, with the most common being Masked Visual-Language Modeling (MVLM) proposed by (Xu et al., 2020). This task is strongly related to the more general Masked Language Modeling (MLM), which was originally introduced with BERT (Devlin et al., 2019). For the sake of MVLM, random tokens of the input sequences are masked where the goal of the model is to predicted the masked tokens given the surrounding context. Importantly, positional information (such as word coordinates in the document image) is not masked and thus helps the model with the reconstruction. In this way, the model learns language contexts while also exploiting positional information. Various other related tasks have been proposed over time, which also explicitly include image modalities. This allows KIE models to fully exploit textual, positional and visual information during pre-training. When authors propose end-to-end and/or generative KIE methods, specific pre-training objectives are defined in order to learn text reading and producing outputs in an autoregressive manner. See also chapter 6.1 for a further discussion.
Regarding the distribution of integrated encoder and decoder architectures, 11 of the 57 sequence-based methods do not incorporate a dedicated textual encoder, while 17 of the methods do not use a visual encoder. In (Palm et al., 2017; He et al., 2023b), neither types of encoders are employed. In most cases, BERT or derived variants such as RoBERTA (Liu et al., 2019b) and SBERT (Reimers and Gurevych, 2019) are used for textual encoding. This is not surprising, given the popularity of BERT-based models in NLP. Another common encoding method is the LayoutLM-family, more specifically LayoutXLM, LayoutLM and LayoutLMv2, used by (Appalaraju et al., 2021; Nguyen et al., 2021; Zhang et al., 2021; Cao et al., 2022a; Du et al., 2022; Gu et al., 2022; Li et al., 2022c; Wang et al., 2022a; Deng et al., 2023a; He et al., 2023c; Li et al., 2023b). This can be explained by the popularity of these models in the context of DU as well as their extensive pre-training, which allows approaches to utilize expressive embeddings. Some authors also make use of more traditional embedding methods such as N-gram-models (Gbada et al., 2023), Word2Vec (Zhang et al., 2021) or FastText (Oral et al., 2020).
Regarding visual encoders, one can see that there is no clear dominance of one particular model. Nonetheless, the most common visual backbone category is represented by ResNet models (He et al., 2016), which are also popular in related CV tasks. ResNet50 and ResNet18 are the most often used variants of this architecture (Yu et al., 2020; Appalaraju et al., 2021; Gu et al., 2021; Li et al., 2021c; Wang et al., 2021b; Cao et al., 2022b; Liao et al., 2023; Yu et al., 2023), however other variants that have been proposed as improvements such as ResNeXt101-FPN and ConvNeXt-FPN are employed as well (Xu et al., 2021; Gu et al., 2022; Dhouib et al., 2023; Luo et al., 2023; Yang et al., 2023a; Yu et al., 2023). Also, (Kim et al., 2022; Davis et al., 2023; Dhouib et al., 2023) incorporate the more recent Swin Transformer (Liu et al., 2021) as the visual encoder, which captures local and global image features through sequential processing of non-overlapping image patches and shifted window self-attention mechanisms. Interestingly, only the approach by (Guo et al., 2024) integrates a Vision Transformer (ViT). This model also leverages transformers for visual tasks, however compared to Swin Transformers, the input images are processed as sequences of patches without hierarchical divisions. Document Image Transformer (DiT) (Li et al., 2022b), which in turn is a self-supervised improvement over ViT, is also only used in the approach by (Huang et al., 2022).
The distribution of chosen decoders shows that the majority of approaches use sequence labeling layers for decoding. This usually consist of a linear and a Softmax layer, which assigns a probability distribution over all possible fields to each token in the sequence. The task of KIE is then performed by choosing the field with the highest probability according to the probability distribution. Besides, some approaches alternatively use LSTMs (Hochreiter and Schmidhuber, 1997), BiLSTMs (Schuster and Paliwal, 1997), Conditional Random Fields (CRF) (Lafferty et al., 2001) or a combination thereof (Palm et al., 2017; Guo et al., 2019; Oral et al., 2020; Sage et al., 2020; Yu et al., 2020; Hamdi et al., 2021; Klaiman and Lehne, 2021; Ning et al., 2021; Wang et al., 2021a, b; Kuang et al., 2023; Zhang et al., 2023a). BiLSTMs are often used as they effectively capture contextual information in a bidirectional fashion. CRFs are useful, as they model the dependencies between labels in a sequence and incorporate a global context. These more complex models can therefore provide additional information compared to a simple Softmax layer and can potentially increase the extraction performance. When looking at the decoder choice over time, one can see that the two aforementioned variants have not been used as frequently since 2022. Also, especially in 2023, newer decoding methods were examined. Some authors also use existing KIE approaches for decoding. For example, (Hong et al., 2022) employ the approach proposed by (Hwang et al., 2021b) as the decoding mechanism and (Deng et al., 2023a) base their work on the system by (Kim et al., 2022). Besides the already mentioned architectures, a plethora of different models are employed. We refer to table LABEL:sequencetab for a listing of the methods.
All in all, it can be stated that there are many different approaches with regard to the employed encoder and decoder architectures. BERT-variants are most often used for textual encoding, while ResNet-based models are used as visual backbones. Decoding steps are usually performed either by simple sequence labeling layers or more complex architectures that allow to incorporate certain dependencies within the token sequences.
5.2.2. Graph-based methods
The findings regarding the analyzed graph-based approaches are presented in table LABEL:graphtab of the electronic supplementary material. A key property in which the approaches differ, is how the underlying graph representation is constructed. 10 of the approaches use a fully connect graph (Liu et al., 2019a; Carbonell et al., 2020; Luo et al., 2020; Tang et al., 2021; Zhang et al., 2021; Wang et al., 2022a; Zhang et al., 2022a, b; Gemelli et al., 2023; Zhang et al., 2023b), which means that every single node of the graph is connected to each other. As a result, the resulting graphs usually have a very large number of edges. In 8 cases, a k-nearest neighbors algorithm is chosen, where each node only has k neighboring nodes. The approaches use different values for k, namely 4 (Zhang et al., 2022b; Belhadj et al., 2023a; Hamri et al., 2023), 5 (De Trogoff et al., 2022), 8 (Belhadj et al., 2023a), 10 (Carbonell et al., 2020), 15 (Deng et al., 2023b) or even 36 (Zhang et al., 2023c). Other common methods are to identify the nearest neighbors in the four major directions up, down, left and right (Lohani et al., 2019; Qian et al., 2019; Gal et al., 2020; Krieger et al., 2021; Li et al., 2023b) or using the ß-skeleton graph algorithm, which connects nodes based on their geometric proximity depending on the parameter ß, where all approaches choose ß=1 (Lee et al., 2021, 2022, 2023). (Holeček, 2021; Wang et al., 2023a) divide documents into distinct areas starting from each node (e.g., in 45 degree angles) and identify closest neighboring nodes in each of these segments. (Cheng et al., 2020) emit 36 rays out of each node and declare all other nodes, whose bounding box is crossed by these rays, as a neighbor. In (Yu et al., 2020; Davis et al., 2021; Hwang et al., 2021b; Gemelli et al., 2023; Wang et al., 2023a), the graph creation is iteratively learned by a neural network and thus not defined by a deterministic algorithm.
Regarding different methods, how the information is propagated through the graphs, in most cases (n=10), a regular Graph Convolutional Network (GCN) is employed (Lohani et al., 2019; Qian et al., 2019; Gal et al., 2020; Luo et al., 2020; Wei et al., 2020; Davis et al., 2021; Holeček, 2021; Lee et al., 2022; Wang et al., 2022a; Lee et al., 2023). GCNs propagate information between the graph nodes to iteratively learn representations, which capture both local and global contexts. Each iteration (i.e., graph convolution) calculates new embeddings for the nodes and edges. In (Liu et al., 2019a; Yu et al., 2020; Tang et al., 2021; Zhang et al., 2021; Deng et al., 2023b), a variation of a GCN, which is defined on node-edge-node-triplets, is used in conjunction with a Multilayer perceptron (MLP). In this setting, the features of the node itself, the edge and all neighboring nodes are fused in order to obtain the new node embedding. Another common method for graph propagation is using Graph Attention Networks (GATs), either with our without multi-head attention (Carbonell et al., 2020; Hua et al., 2020; Belhadj et al., 2021; Krieger et al., 2021; De Trogoff et al., 2022; Zhang et al., 2022a, b; Belhadj et al., 2023a, b; Zhang et al., 2023c). This architecture utilizes the attention mechanism based on transformers to calculate different weight coefficients for each neighboring node, allowing the model to focus on the most relevant information during graph propagation and thus improving the overall node representations (Veličković et al., 2018). For further details regarding propagation types, we refer to the listing in table LABEL:graphtab.
With respect to the graph propagation, one can also analyze the number of propagation layers. The number of layers not only determines the overall complexity of the model, but also the receptive field of each node, determining how far it can aggregate information from its neighboring nodes. Each layer increases the depth of this receptive field by one. On average, the approaches consist of 3.6 layers. (Liu et al., 2019a) conducted ablation studies regarding different numbers of layers and found that the optimal number of GCN layers is 2, which however can be task-dependent and should be considered on a case-by-case basis. With regard to the GAT-based approaches, one need to decide on the number of attention heads. (Belhadj et al., 2021) conducted a study with different attention heads and came to the conclusion that in their case, 26 attention heads produced the best results. The authors also mention that correlations between the number of attention heads and the number of fields to be extracted might exist. Here, too, it is necessary to examine which variant works best in each individual case.
The approaches use different methods how KIE is performed, i.e., how the graph structure is used to extract information from documents. Most of the approaches use a simple node classification (Lohani et al., 2019; Gal et al., 2020; Belhadj et al., 2021; Davis et al., 2021; Krieger et al., 2021; Lee et al., 2021; Belhadj et al., 2023a, b; Hamri et al., 2023), i.e., after the embeddings of each graph node have been obtained, they are used to classify each node into one of the fields to be extracted. Also common are BiLSTMs, CRFs, their combination in the form of BiLSTM-CRFs (Liu et al., 2019a; Qian et al., 2019; Hua et al., 2020; Yu et al., 2020; Tang et al., 2021; De Trogoff et al., 2022; Zhang et al., 2022b; Deng et al., 2023b) as well as sequence labeling layers (Wei et al., 2020; Wang et al., 2022a; Li et al., 2023b; Zhang et al., 2023c). Besides, multiple other methods are used to perform the KIE task. For example, (Lee et al., 2022) use the Viterbi algorithm, which identifies the most likely sequence of states by iteratively calculating probabilities and backtracking through the graph.
In by far the most cases (n=18), the graphs are defined on word-level, which means that every word on a document image represent one node in the constructed graph. This is also inline with the granularities for the sequence-based methods. Another common way to define graphs is to consider entire segments or sentences (n=13). The resulting graphs are therefore rather coarsely defined. 4 of the approaches consider entire text lines as nodes, regardless of whether the contained text elements are related to each other or not (Davis et al., 2021; Tang et al., 2021; Zhang et al., 2022b; Li et al., 2023b). (Wang et al., 2022a) and (Zhang et al., 2022b) define the graphs on two different granularities simultaneously, with one being more fine-granular (either words or even individual token) and the other being more coarse-granular (text segments or text lines).
Another central component of graph-based methods is the definition of node features. These features determine, which types of information the approach can utilize for the KIE task. The analysis has shown that various different features have been proposed over time. In general, the node features usually consist of position-oriented values such as normalized coordinates of the node within the document image. In (Lohani et al., 2019; Luo et al., 2020; Hwang et al., 2021b), relative distances to neighboring nodes are also part of the node features. Another common feature type is related to the textual content represented by the node, e.g., word embeddings. The approaches differ in terms of which embeddings models are used. In this regard, BERT-based-embeddings are often employed, which was also the observation regarding textual encoders of sequence-based approaches. Some other methods are Byte Pair Encoding (Sennrich et al., 2016), BiLSTM and LayoutLM. In some cases, multiple text embeddings are used simultaneously, for example (Zhang et al., 2021) employ Word2Vec, BERT and LayoutLM for the word embeddings. On top of positional and textual embeddings, many approaches also integrate visual features, often obtained from models such as ResNet or LayoutLM and their variants. Besides the previously mentioned features, many authors also use hand-crafted miscellaneous features. The goal of these features is usually to incorporate certain node characteristics as well as relationships between nodes that should aid the model in performing KIE. (Lohani et al., 2019) for example include various boolean features whether a graph node represents a date, zipcode or known city, among others.
The approaches also differ regarding the size of the node feature vectors. Although in many cases the actual number is not reported in the manuscripts, on average, the features vectors have a length of around 550. This is somewhat smaller compared to common embeddings sizes such as 768, which are used by the original BERT model (Devlin et al., 2019). In general, there is little discussion about embeddings sizes for node features in related work, from which it can be concluded that these have no significant influence on the performance of the KIE systems. Much more important seem to be the different types of features which are being integrated. Another distinguishing factor of graph-based approaches is, whether they incorporate a global node in the defined graph representation. This global node can contain information about the entire document image and can be connected to the individual nodes of the graph. To this end, only the approaches by (Hua et al., 2020; Zhang et al., 2023c) use such global nodes. (Li et al., 2023b) also construct a global node, however it is not used as a information carrier and is rather formally required due to the chosen concept based on document layout trees.
Edges represent the second integral part of the graph structures. It is noticeable that a large portion of work (n=17) does not incorporate any edge features. Consequently, less attention is paid to this type of feature compared to node embeddings. In the other cases, edge features are usually derived from pairwise positional relationships between the connected nodes. Most often, horizontal and/or vertical distances, both in absolute and relative values, are considered (Liu et al., 2019a; Wei et al., 2020; Tang et al., 2021; Zhang et al., 2021). The authors highlight the importance of the visual distances between two nodes (Liu et al., 2019a). The aspect ratio of the bounding boxes that span the graph nodes is also often included as edge features (Liu et al., 2019a; Yu et al., 2020; Tang et al., 2021; Lee et al., 2022, 2023). In some papers, certain edge types are defined as edge features. For example, (Qian et al., 2019) use the the edge direction (left-to-right, right-to-left, up-to-down, down-to-up) as one feature value. In only three of the graph-based approaches are visual embeddings part of the edge features (Davis et al., 2021; Lee et al., 2021, 2023). Regarding the distribution of edge directions, i.e., whether the edges are directed or undirected, there is no clear dominant variant, however in the majority of the cases (n=19), the edges are undirected. (Tang et al., 2021; Zhang et al., 2021, 2022a) did not explicitly state which type they use and it is also not clear from the manuscript. In general, the choice of the edge definition does not seem to have a significant influence on the performance of the corresponding KIE systems. Additional research is required to appropriately estimate the impact of certain design features such as the edge direction on the extraction performance.
5.2.3. Grid-based methods
As mentioned in section 5.1, this paradigm is less popular in related literature and only 10 approaches make use of such grid structures. The analysis of these is shown in table LABEL:gridtab of the electronic supplementary material. An interesting observation is that most of the identified approaches represent evolutions of existing grid-based systems, in particular based on Chargrid (Katti et al., 2018). The only methods that propose novel concepts are (Katti et al., 2018; Gal et al., 2019; Palm et al., 2019; Wang et al., 2021b). This shows that related research is mostly conducted within a very narrow corridor. Regarding the chosen granularities, a majority of the systems is defined on word level (Denk and Reisswig, 2019; Gal et al., 2019; Palm et al., 2019; Kerroumi et al., 2021; Lin et al., 2021; Do et al., 2023), while three are defined on character level (Katti et al., 2018; Dang and Thanh, 2020; Yeghiazaryan et al., 2022) and one method on token level (Wang et al., 2021b).
Table LABEL:gridtab also lists findings regarding the grid dimensions. Almost all methods define the height and width of the grid based on the pixel counts of the input image. The grids are therefore rectangular, as this corresponds to the rectangular format of corresponding VRDs. (Gal et al., 2019) and (Yeghiazaryan et al., 2022) use a fixed value for the height and width, namely 512 and 336 respectively, therefore resulting in square grids. (Kerroumi et al., 2021) scale down the input dimensions by a factor of 8 in order to reduce the overall model complexity. The work of (Wang et al., 2021b) use a somewhat different approach and employ a dynamic grid structure, in which the height and width dimensions are defined by the distance between the maximum and minimal coordinates of bounding boxes in vertical and horizontal directions, i.e., the span of actual document content. Another distinguishing factor is the feature vector length. Relatively similar values are chosen here, with 256 and 768 being the most common vector sizes. The approach of (Gal et al., 2019) uses the smallest feature set with a vector of length 35. The most complex feature vector is defined by (Palm et al., 2019). In this case, the feature vector has the dimension 4x128x103.
Regarding the design choice, which features are chosen for the individual grid elements, the majority of approaches use text embeddings obtained from different models. A common choice are BERT-based models including BERT, RoBERTa and PhoBERT (Denk and Reisswig, 2019; Lin et al., 2021; Yeghiazaryan et al., 2022; Do et al., 2023). Another common feature type is a 1-hot encoding of the characters (Katti et al., 2018; Palm et al., 2019; Dang and Thanh, 2020), which means that each character in the word corpus is assigned to a specific index value. Only the approaches by (Gal et al., 2019; Lin et al., 2021; Wang et al., 2021b; Do et al., 2023) integrate visual features, which are based on pixel-wise RGB-values, ResNet18, or Swin Transformer embeddings. Another distinguishing feature is the integration of features for background elements of the grid, i.e., all elements that are not overlapped by contents of the document image. In this regard, almost all approaches use an all-zero feature vector. The aim is to clearly differentiate actual content and backgrounds. (Palm et al., 2019) use a sparse tensor and therefore discard background elements entirely. Only the work of (Kerroumi et al., 2021) integrate features for the background elements, namely the RGB-channels of the corresponding coordinates in order to fuse visual information into the architecture.
Most of the methods employ a semantic segmentation (Katti et al., 2018; Denk and Reisswig, 2019; Dang and Thanh, 2020; Kerroumi et al., 2021; Lin et al., 2021; Yeghiazaryan et al., 2022) to perform the KIE task, where each element of the document image is assigned to semantic categories (i.e., fields to be extracted), which results in a segmentation mask of the original document image. Other choices are bounding box regression (Katti et al., 2018; Denk and Reisswig, 2019), where the goal is to predict bounding boxes of semantic entities, which can be helpful to better differentiate individual objects. For example, when processing an invoice, it is helpful to separate each individual line item. This is also the reason why (Katti et al., 2018; Denk and Reisswig, 2019; Yeghiazaryan et al., 2022) combine semantic segmentation with methods for bounding box regression (or line item detection in general). We refer to table LABEL:gridtab for a detailed overview regarding employed KIE methods.
5.2.4. Other methods
Of the 96 papers analyzed, 6 cannot be assigned to one of the paradigms of KIE methods discussed before. In the following, we briefly present key concepts behind those deviating approaches. The approach by (Majumder et al., 2020) is based on learning representations document snippets and classifying them into fields to be extracted. First, potential candidates are identified for each field based on the data type. The next step is to select the correct candidate for each field. To this end, a representation is obtained for each candidate as well as each field to be extracted. The candidate representation is constructed based on the candidate itself as well as its neighboring segments, utilizing textual and positional information processed by a self-attention mechanism. Finally, a similarity score is calculated for each pair of candidate embedding and field embedding. The similarity scores are then used in a separate module to select the appropriate candidate for each entity, which can be defined in a variety of ways. A trivial method is to select the candidate with the highest similarity score for each entity.
(Wang et al., 2020) use a hierarchical tree-like structure to represent fragments of the document pages, which can be compared to the graph-based methods. Compared to traditional tree architectures, nodes on the same hierarchy level can also be connected with each other. Parent and child nodes of these trees represent key-value pairs. The KIE task is performed by predicting the relations between the individual fragments, i.e., directions of the edges. Obtaining the most likely parent element of each fragment can be ultimately used for identifying fields to be extracted.
(Zhang et al., 2020) propose an end-to-end KIE system and combine text reading and text understanding. The text reading steps includes a detector model as well as an LSTM-based recognition model which identify text in the document images. The approach also includes a dedicated module to fuse the different input modalities, which is then used by the final KIE module. A BiLSTM coupled with a fully connected network is employed as a decoding mechanism to predict relevant entities for the fields of interest. This approach is therefore closely related to sequence-based KIE methods.
(Sarkhel and Nandi, 2021) follow concepts of (Majumder et al., 2020) in the sense that for each field to be extracted, a set of candidate spans is identified first. In order to identify candidate spans for a named entity, multiple detector functions are implemented, to some extent based on domain-knowledge. The authors then employ an adversarial neural network in order to find local contexts of the identified visual spans. These local contexts represent specific fragments of a document image that contain the spans as well as related relevant context (e.g., a larger paragraph). Based on these identified segments, both global and local context vectors are constructed based on textual and visual features. These features are then used by a binary classification model in order to predict, whether a visual span contains a specific field to be extracted or not.
The approach by (Tata et al., 2021) is another method that is based on generating candidates for each field and scoring them subsequently. Given a document image and fields to be extracted including their data types, the system first identifies candidates in the OCR-output based on third-party detector functions. The identified candidates are then scored according to their likelihood of correctly representing a particular field of interest in a binary classification setting. Similar to (Majumder et al., 2020), the score depends on the similarity between the embeddings of the candidate and the field of interest. Different scoring functions can be integrated, however the authors choose a function that assigns the candidate with the highest similarity score to each field.
(Shi et al., 2023) propose a novel document representation structure which they refer to as cell-based. This representation is similar to grid-based methods, however the cell-based methodology has no consistent placement of elements in terms of height and width. Instead, the cells are defined depending on the actual document content. For example, different lines or columns in the cell-structure can consist of a different number of elements. The individual cells are also sorted by row and column index respectively, which provides additional information to the KIE system. The obtained cell-based layout is then processed by sequence-based methods such as LayoutLM and therefore follow the typical sequence labeling scenario that respective models use to perform KIE.
5.3. Evaluation
5.3.1. Methodology and setup
Table LABEL:evaltab of the electronic supplementary material shows the findings regarding the presented evaluation setups. In 18 of the analyzed papers, an element-based evaluation method is chosen, which means that elements like the predicted class of a token are compared with the groundtruth class in case of sequence-based approaches. Another common element-based evaluation is to compare the predicted bounding box with the actual bounding box and to identify a match or mismatch based on the overlap of their respective coordinates. 32 of the analyzed approaches incorporate a string-based evaluation setting. In such cases, the extracted textual values are compared with the groundtruth strings to determine, whether a prediction for a given field was correct or not. A string-based evaluation is particularly relevant for assessing the suitability for real-world applications where the extracted texts are used for further processing (e.g., transfer to other information systems). In cases where the authors either did not explicitly specify their chosen method and/or it was not obvious from the manuscript itself, we declared the property as ”n/a”. This was the case for half of the analyzed papers, showing that authors often do not describe their evaluation procedures in great detail.
In almost all of the manuscripts, the authors present the overall performance, i.e., the aggregated results across all fields of interest and all documents of the test set. However, the field-level performance, which shows the performance for each field individually, is only adopted in 35 of the analyzed papers. One aspect that is almost never presented is the performance on the Unknown class. In the context of KIE, Unknowns represent all elements of a document that do not belong to any of the fields to be extracted. This can be helpful in determining the extent to which an approach is able to distinguish irrelevant content from relevant parts of a document image. However, only (Krieger et al., 2021; De Trogoff et al., 2022; Gemelli et al., 2023) explicitly report the evaluation results for this class. In most cases (n=85), the proposed method is compared against existing KIE systems, either by comparing the own results with the evaluation metrics presented in the respective papers or by re-implementation. Thus, it can be stated that a high emphasis is placed on the overall comparability and the legitimization of the own approach. However, especially when authors decide to re-implement existing approaches, comparisons of evaluation results are not always appropriate, since less effort is usually put into training the approaches of third parties.
Considering the employed evaluation metrics, one can see that the F1 score is predominantly used, more specifically in 81 manuscripts. Precision and Recall, on which the F1 score is based, are only explicitly reported in about 25 % of the cases. Interestingly, the usually widely used Accuracy metric is only adopted by (Gal et al., 2019; Palm et al., 2019; Cheng et al., 2020; Gal et al., 2020; Belhadj et al., 2021; Klaiman and Lehne, 2021; Ning et al., 2021; Sarkhel and Nandi, 2021; De Trogoff et al., 2022). One reason for this is that the task of KIE is a multi-class classification problem and that the datasets are usually very unbalanced in terms of the distribution of the fields to be extracted compared to the Unknowns, which strongly distorts the Accuracy metric (Galar et al., 2012). Moreover, it is noticeable that a large majority of the identified evaluation metrics (15 out of 24) are custom-defined and only used in one particular paper, which prevents an adequate benchmark against existing KIE approaches.
Custom datasets are being used in 45 of the analyzed papers and typically represent in-house document collections that are not publicly available. However, in most of the cases when custom datasets are used, the authors also present their results on public benchmark datasets to allow for a comparative evaluation. In addition, there are numerous datasets that were proposed in individual papers, but were not adopted in any other related work. This is the case for 16 of the 26 identified datasets. On average, two datasets are used for evaluation purposes.
Table 2. Common datasets for KIE research
Dataset Documents # docs # classes555in terms of entities to extract Language # uses666as part of analyzed approaches Pre-Training?
FUNSD Forms 199 4 eng 43
CORD Receipts 1,000 30 eng 34
SROIE Receipts 973 4 eng 32
IIT-CDIP Lawsuits 6,000,000 / eng 21
EPHOIE Exams 1,494 10 zho 7
RVL-CDIP Lawsuits 400,000 / eng 6
Kleister-NDA NDAs 540 4 eng 5
XFUND Forms 1,393 5 zho,jpn,spa,fra,ita,deu,por 5
DocBank Miscellaneous 500,000 12 eng 4
Table 2 provides an overview of the most common datasets adopted in KIE research. Listed are benchmark datasets as well as pre-training datasets. By far the most common datasets in this regard are FUNSD (Jaume et al., 2019), CORD (Park et al., 2019) and SROIE (Huang et al., 2019), all of which are used by at least a third of the papers. FUNSD includes forms from various domains. CORD and SROIE on the other hand contain photographed receipts, mostly from supermarkets and restaurants. There are discrepancies both in terms of dataset size and in the number of fields/keys to be extracted. The datasets with many hundreds of thousands of documents are typically used for the pre-training of corresponding models and not as an evaluation benchmark. Therefore, these datasets are also widely adopted in other DU tasks. Two of the three most used datasets only aim at extracting four fields of interest, which is a relatively low amount compared to the variety of information corresponding documents usually include. CORD on the other hand includes 30 fields to be extracted, which is also the highest among the listed datasets. The key difference between CORD and SROIE is that in case of former, detailed information including individual line item attributes such as their quantity or unit price need to be extracted, while SROIE only requires to extract aggregated information such as the total price. For the most part, the datasets include English documents. One exception is XFUND (Xu et al., 2022), which specifically intends to investigate multi-lingual capabilities of KIE approaches. Two datasets based on Chinese documents exist, namely EPHOIE (Wang et al., 2021a) and Ticket (Guo et al., 2019), however they are understandably not as widely adopted as most English counterparts that are being used by the international research community.
5.3.2. Quantitative comparison
We now discuss the evaluation results for the common benchmark datasets CORD, FUNSD and SROIE. We only consider the F1 score as it represents the most common evaluation metric. Note that we do not distinguish between micro, macro or weighted F1 averages [^7] For an explanation of these aggregated metrics, we refer to (Sokolova and Lapalme, 2009)., since in many cases it has not been explicitly specified by the authors. The results should therefore be treated with caution, especially since macro averages usually tend to be lower than the micro averages.
To this end, figure 6(c) shows the results for the three datasets given all papers that present results on the corresponding datasets. A trend line is also displayed. In all cases, one can see that there is no steep increase in performance over time. However, the largest improvement over time was achieved with FUNSD, as can be seen from the trend line. In the case of SROIE, the trend line is almost horizontal, therefore indicating a stagnation. All graphs show a few outliers that achieved worse results than previous and subsequent work. One reason for this may be a different calculation of the F1 scores, e.g., by determining macro average values instead of micro average ones. The best performing approach for CORD is #28 proposed by (Gu et al., 2021), for FUNSD #85 by (Li et al., 2023b) and regarding SROIE #96 by (Zhang et al., 2023c). The first two are sequence-based approaches, while the latter is a graph-based method. On average, very comparable results are obtained for CORD and SROIE, which can be attributed to the fact that they both contain receipts. The average F1 score for both datasets is 0.96. In case of FUNSD, however, the results are significantly worse, with an average F1 score of only 0.82. One problem that KIE approaches face with FUNSD may be that this dataset was not primarily constructed for the KIE task. The fields to be extracted often span multiple lines of text, which is very different from traditional KIE datasets that aim at extracting key information such as a specific date.
(a) CORD
(b) FUNSD
(c) SROIE
Figure 6. Evaluation results and correlation between model size and performance
[…]
Graphs that show the performance on three benchmark datasets and the fact that there not necessarily exists a correlation between model size and performance.
Figure 6(c) visualizes the relation between model sizes in terms of parameter count and performance in terms of F1 score for the three datasets. Note the logarithmic scale of the x-axis due to the large range of possible values. In all cases, there is a clear cluster formation with some outliers. The correlation coefficient for CORD is -0,10 (p=0,62), for FUNSD 0,13 (p=0,49) and for SROIE 0,12 (p=0,65). Therefore, there is no significant correlation (neither negative nor positive) between model size and performance. Interestingly, the correlation in case of CORD is even negative, which means that many of the approaches with smaller parameter counts are able to outperform more complex KIE methods. The figure also differentiates the KIE paradigms by color. One interesting observation is that regarding FUNSD, especially graph-based methods often produce worse results compared to other paradigms and generally represent outliers. One reason for this could be the special layout of the forms in this dataset. As mentioned before, the forms typically contain entities that span multiple text lines. A constructed graph, independent of the chosen granularity, is indisputable an unfavorable document representation for such layouts. At the same time, one can see across all datasets (especially FUNSD), that a combination of graph-based and sequence-based modules seems to increase extraction performance compared to a purely graph-based approach. The figure indicate that increasing the model size beyond a certain threshold does not necessarily correlate with a significant increase in performance and that creating more complex models can even yield diminishing performance returns. The fact that all paradigms can achieve competitive results (with the exception of graph-based methods w.r.t FUNSD) also suggests that the choice of model architecture can be flexible depending on the specific application. Therefore, approach design seem to be more crucial than the sheer parameter count. This is a relevant insight especially for practical applications, where resources and computational cost need to be taken into account.
In case of FUNSD, the spread is much larger compared to the other two datasets. Here, many of the approaches with fewer parameters achieve significantly worse results compared to rest. One can conclude that the challenging FUNSD dataset with its unique entities therefore represents a particular challenge for KIE systems of lower complexity. The approach #81 proposed by (He et al., 2023b) with 175 billion parameters is not able to outperform most of the other approaches with significantly fewer parameters. This is especially true for CORD, where it is one of the worst performing approaches. However, competitive results are achieved for the other datasets, especially SROIE. Approach #75 by (Dhouib et al., 2023) with 70 million parameters represents another outlier and achieves less competitive results regarding CORD and SROIE, even though the model size is not significantly smaller compared to the other methods. The system by (Hamri et al., 2023) only contains 53,6 thousand parameters and is therefore the by far smallest one compared to the other analyzed KIE methods. Nonetheless, the approach can achieve convincing results in case of SROIE and even outperforms the aforementioned method with 70 million parameters.
6. Discussion
6.1. Pre-Training and Fine-Tuning procedures
An increasingly popular approach in KIE research, in particular when dealing with sequence-based methods, is to split model training into self-supervised pre-training and supervised fine-tuning steps. The aim of the former is to acquire a general document understanding by the model first, which can then be utilized for different downstream tasks such as KIE as part of subsequent fine-tuning steps. In general, pre-training procedures require large volumes of data to properly let models learn corresponding document representations, which is why most authors use large datasets such as IIT-CDIP, as also discussed in chapter 5.2.1. The problem with this dataset in particular is that included documents originate from the 1990s and therefore show poor image quality, noise produced by faulty scans as well as a generally low image resolution. This also means that a large proportion of the proposed approaches are pre-trained on outdated documents and may therefore have difficulties when they are fine-tuned on newer documents as part of real-world use cases. Another aspect that should be considered is the fact that all documents of IIT-CDIP are related to lawsuits against the tobacco industry and therefore lead to a strong domain bias that is being adopted by the models during pre-training. In this respect, it is debatable whether it would not be more effective if pre-training procedures were carried out based on more representative datasets (see also chapter 7).
As discussed in chapter 5.2.1, the majority of authors adopt MVLM during pre-training. In some cases, it is even the only pre-training objective that is being used. Given the general popularity of MLM in the NLP domain, it is not surprising that the MVLM task, which is derived from it, is also popular in KIE research. It represents an effective method to let models learn useful representations given the surrounding textual and positional contexts. As the adequate processing of VRDs also requires to consider visual cues, newer pre-training objectives have been proposed in order to fuse visual information into KIE systems. This can be done in different ways, for example by letting the model reconstruct local or global document image snippets (Appalaraju et al., 2021; Tang et al., 2023) or predicting the segment length of an image snippet (Li et al., 2021c). Other possibilities aim at various matching-tasks between text and document images, for example where the model has to predict whether a sentence describes a document image (Appalaraju et al., 2021), if a given image and text are part of the same page (Xu et al., 2021) or if image patches of a word are masked or not (Huang et al., 2022). Some authors also define additional pre-training objectives to better exploit positional information, for example by the means of classification problems to estimate a tokens placement into a document image area (Li et al., 2021a; Wang et al., 2022b) or to estimate relative positional directions of an element to its neighbors (Li et al., 2021c). Research is also conducted related to a better understanding of numerical values and their relationships in business documents (Douzon et al., 2022). This can be beneficial for processing document types such as invoices, which usually contain various monetary values that are closely related to each other. To conclude, the proposed pre-training objectives are carefully designed to allow for an adequate fusion of textual, positional and visual inputs, which helps corresponding models to obtain a document understanding which can then be exploited in downstream tasks like KIE.
Another category of pre-training objectives stems from the generative KIE methods, where corresponding models need to learn both natural language understanding and natural language generation (i.e., generating output sequences conditioned on input sequences) capabilities. To this end, (Cao et al., 2022a, b) adopt the pre-training tasks defined by (Dong et al., 2019), namely Unidirectional LM, Bidirectional LM and Sequence-to-Sequence LM. Importantly, all pre-training tasks are considered simultaneously by using a shared transformer network which can alternate between the three objectives. (Tang et al., 2023) propose a range of self-supervised and supervised generative pre-training objectives based on task prompts and target outputs in textual form, for example where the model must predicts missing texts and locate them in document images as a structured target sequence. The approach by (Davis et al., 2023) follows a specific strategy for learning text recognition, document understanding and generative capabilities. It generally stands out in terms of its pre-training setup, as over 25 different objectives across different document types are being used. Nonetheless, a strong focus is placed on MLM-related tasks. One important pre-training task revolves around the model predicting a structured JSON output for a given document, which is also ultimately used for KIE. This strategy can be helpful in real-world settings, in which a detailed hierarchical output of input documents is required, such as an adequate differentiation between individual line items in invoices. Non-generative KIE methods that for example decode outputs with a sequence labeling layer usually have difficulties in reconstructing such hierarchies and require additional post-processing steps. To let the model learn text recognition, (Kim et al., 2022; Dhouib et al., 2023) use a pre-training task where the model must predict the next token while considering the previous tokens as well as the document image.
6.2. Practical perspective
Based on the analysis, there is a clear lack of a practice-oriented perspective that is being adopted in related work. To this end, only a small number of the analyzed approaches integrate domain knowledge. Moreover, in most of these cases, domain knowledge is at most fused into the systems by the means of hand-crafted input features. There is also a lack of evaluating proposed KIE approaches and their impact on real-world scenarios. These observations have also been made in the study by (Martínez-Rojas et al., 2023).
One work that stands out in terms of its consideration of domain knowledge is the approach by (Arroyo et al., 2022). The authors make use of such knowledge by proposing a hybrid KIE system based on both DL-models and rule-based methods including several post-processing steps to improve the extraction results. This includes, for example, the correction of automatically extracted product codes and product prices. However, this post-processing is considered as a supplement to the DL model in case of incorrect predictions and not an architecture-based adaptation. Another example is the work by (Palm et al., 2019), which integrates domain-oriented constraints regarding invoices that the total amount is the sum of the subtotal and tax total values, and that the tax total can be computed as the product of subtotal and tax percentage. The authors achieve this by adjusting the loss calculations during training, but report that no significant performance improvements were achieved. At the same time, this emphasizes the need for further research in this area (see also chapter 7). While not directly implementing such aspects in the proposed methods, some authors also designed their approaches in a way to allow for integrating domain knowledge. For example, in the work of (Tata et al., 2021), candidates are first identified for fields to be extracted and subsequently scored. In this regard, one example the authors mention is to define a scoring function that integrate constraints such as the fact that an invoice date must precede its due date.
As discussed in section 2.1, adopting a practical perspective and considering domain knowledge can provide valuable insights for an automated extraction system in various ways. However, this aspect has not been properly explored in the literature, even though it has been named as a key challenge for DU in the past (Motahari et al., 2021). Therefore, future work should consider adopting a business process perspective in order to develop approaches that explicitly exploit this unrealized potential. This could also lead to the identification of completely new potentials of KIE systems that are not yet considered in current research, as the current focus often lies only on increasing the performance with respect to established benchmark datasets. These considerations may require a deeper understanding of the associated business processes, which can be achieved by modeling computer-integrated systems including data, behavior and control flows (Fettke and Reisig, 2021). The potential of context-aware DL systems has been investigated in other research areas. For example, (Weinzierl et al., 2019) propose the fusion of information extracted from business documents with traditional event log data in order to enhance predictive process monitoring capabilities. However, the other direction, i.e., feeding context-aware data obtained from process mining techniques into KIE systems, has not yet been considered.
6.3. Trends
One can observe several trends in KIE research since 2017 in a wide variety of facets, some of which were also addressed in chapter 5. One aspect is the dominance of the three main paradigms of KIE systems and most notably sequence-based methods, which represent the predominant approach since 2021. To the contrary, grid-based methods have never been able to establish themselves, despite their comparability with graph-based approaches, which are even more popular now (especially in 2023) than they were in the beginning. At the same time, this indicates that VRDs with usually complex layouts cannot be appropriately represented as grids with well-defined structures that are less flexible compared to dynamic graphs. Also, using conjunction of different paradigms (e.g., sequence-based and graph-based) is not as widely adopted as the focus on one particular paradigm.
Since 2022, an increased focus has been placed on creating OCR-independent and/or generative KIE approaches. The latter may be the consequence of the increasing popularity of generative AI methods such as LLMs in recent years and their influence on the DU domain. The increasing shift towards OCR-independent systems can be explained by the otherwise necessary external OCR engines, which can be error-prone and therefore lead to falsely extracted information, especially in real-world use cases (Cui et al., 2021).
It can also be observed that over time, there have been different strategies on how to develop and improve KIE systems. While towards the beginning, a lot of research effort has been placed on how to properly integrate the different input modalities, other strategies have also emerged in recent years. For example, some publications highlight and investigate the importance of the reading order. In traditional methods, and also as a consequence of the reliance on external OCR engines, a simple top-bottom and left-right reading order is usually adopted. This however can lead to a non-ideal segmentation of complex VRDs. To this end, approaches such as those by (Peng et al., 2022; Li et al., 2023b; Shi et al., 2023; Zhang et al., 2023b) investigate more sophisticated methods to obtain an optimized reading order that better suits the actual document layout. However, what was not or hardly ever considered as a lever for better KIE performance, is the model size in terms of trainable parameters. Between 2017 and 2023, there was no significant increase in model sizes, as also discussed in connection with figure 5.
It is also positive to note that code and/or model weights are being published more frequently in recent years. Even if the absolute number of implementations that have been shared is relatively low, it has become increasingly more frequent, especially since 2021. This is a positive development, as it can accelerate research progress and research dissemination. This observation also goes hand in hand with the increased popularity of HuggingFace [^8] https://huggingface.co/, which is a platform that provides tools, datasets, and pre-trained models to facilitate research in NLP and CV. Some previously discussed KIE approaches are also available, for example LayoutLMv3999https://huggingface.co/microsoft/layoutlmv3-base.
It can be assumed that these trends and observations will continue in future KIE research. Especially a shift towards generative KIE systems seems evident, as corresponding models become more and more popular across multiple domains. Initial approaches that make use of powerful LLMs exist, however they do not consistently achieve competitive results compared to specialized DU methods (such as LayoutLM) yet. In this regard, additional research is required on how to close this performance gap and thus make corresponding approaches more viable, especially with regard to the trade-off between model size (and therefore hardware requirements) and extraction performance as discussed in section 5.3.2.
7. Research agenda
We have identified several aspects that should be considered in future work, which could not only lead to better research results, but also to a better applicability of KIE systems in real-world scenarios.
  1. (1)
    Novel datasets: Only a small number of public datasets are commonly used. For example, more than half of the sequence-based approaches are pre-trained using (subsets of) IIT-CDIP. One problem with this dataset in particular is the relatively low image quality, which no longer meets today’s standards. These properties have a direct impact on the use of corresponding approaches in real-world settings, which typically include documents with better image quality.
    In addition to pre-training datasets, efforts should also be made to construct novel benchmark datasets, ideally based on different document types compared to receipts or forms which are most commonly considered, as also discussed in (Li et al., 2023a; Skalický et al., 2022). Besides different domains, newly created datasets should also be designed to more closely resemble documents found in real-world scenarios. (Yang et al., 2023a; Zhang et al., 2023b; Wang et al., 2023b) have shown, that existing dominant datasets have numerous shortcomings in this respect. Also, common benchmarks such as SROIE and FUNSD show a high degree of layout replication in training and test partitions (Laatiri et al., 2023), which distorts the overall validity the reported evaluation results. One possible solution to construct novel datasets can be the generation of synthetic documents, however, there is a risk that synthetic documents do not have the properties of real-world documents in terms of layout variety and complexity (Skalický et al., 2022).
  2. (2)
    Consistent evaluation: The analysis in section 5.3 has shown that the papers are very heterogeneous in terms of their evaluation setup. In addition, the authors often do not specify their evaluation methodology in much detail, which raises questions on how exactly the evaluation results were obtained. Therefore, a quantitative comparison of the approaches is not always meaningful. This is especially problematic since the understanding of improving the state of the art in this research area is often associated with an increase in performance on benchmark datasets compared to existing work. Some public competitions, such as SROIE, provide a dedicated evaluation protocol for a consistent ranking of methods.  [^10] The leaderboard for the KIE task can be found at https://rrc.cvc.uab.es/?ch=13&com=evaluation&task=3. However, this is the exception rather than the rule.
    Therefore, it would be desirable to agree on a consistent approach or implement a centralized evaluation toolkit that could be used for different benchmark datasets, where a specific set of metrics (e.g., Precision, Recall, F1) is implemented. (Borchmann et al., 2021) have proposed a reference implementation for a DU benchmark, but it is not yet widely used.
    We also advocate a more frequent use of string-based evaluation metrics, as they allow for a better assessment in real-world applications. However, only one third of the analyzed manuscripts used such string-based evaluation setups. It is also surprising that only about a third of the papers present the results on field-level. In this regard, we advocate that such an evaluation should be presented more frequently in future work, as it gives a good indication of whether an approach faces problems with certain fields, which in turn allows for a more in-depth analysis. Last, but not least, authors should adopt entity-level performance measures, as it is more suitable to assess the performance for practical applications where the extracted entities are processed as a whole (Zhang et al., 2023b).
  3. (3)
    Pre-processing methods: There exists a dedicated research area that investigates techniques for pre-processing document images. An example of a sophisticated approach is proposed by (Saifullah et al., 2023b). In this context, synergy effects with this research area should be exploited, in particular the integration of corresponding techniques into KIE systems, as also discussed by (Baviskar et al., 2021). Related work often does not include pre-processing techniques, or, if they are used, only simple methods such as Deskewing. During the life cycle of a document in real-world settings, there are various points at which image quality can be impaired, for example due to introduced scan artifacts by digitization steps or as a consequence of multiple compression and transmission procedures (Alaei et al., 2024). (Yang et al., 2024) show, that quality aspects such as image noise or fonts can have significant impacts on the performance of DU systems. Therefore, more emphasis should be put on the adoption of respective methods in DU pipelines.
  4. (4)
    Tokenization: In particular sequence-based KIE approaches utilize tokenizers for the (textual) encoding of input documents into token sequences, which are subsequently supplemented by visual and positional embeddings. Whenever the KIE systems incorporate pre-trained language models such as BERT as their encoder backbone, they usually also adopt the corresponding tokenizer without further adjustments. However, there have been studies highlighting the challenges associated with the adoption of unaltered tokenizers to different domains (Nayak et al., 2020). To this end, several methods have been proposed to adjust tokenizers to new domains, which improves the performance on downstream tasks. Examples are domain-specific augmentations of the original tokenizers’ vocabulary (Tai et al., 2020; Sachidananda et al., 2021), or a careful investigation of training data, pre-tokenization setups and other changes to the vocabulary (Dagan et al., 2024).
    Currently, there is no extensive investigation of the role of tokenizers in the context of KIE. (Theodoropoulos and Moens, 2023) analyze the impact of tokenization regarding NER from biomedical texts, however there is a lack of research considering complex DU tasks on VRDs. Therefore, future work should consider exploring sophisticated methods to allow for an adequate transfer of existing tokenizers to KIE and/or develop novel methods for tokenizing complex documents in a more robust manner. Conceivable first steps could be to initially use existing KIE approaches based on out-of-the-box tokenizers, only focus on the adaptation of the tokenizer and perform benchmarks against the original implementation.
  5. (5)
    Generalizability: (He et al., 2023a) have shown that existing KIE methods lack adequate generalization capabilities, which may be due to the limitations associated with the datasets as mentioned before. Furthermore, the aspect of generalizability is often not considered in detail. Instead, it is usually only implicitly considered when the approaches are evaluated based on multiple benchmark datasets with different document types.
    Future work should focus on how to effectively design model architectures as well as novel (pre-)training tasks in order to achieve a higher degree of generalizability across different document layouts, types and domains. Another possibility is to investigate the modularization of KIE systems more closely in order to reduce interdependencies between individual components and thus obtain more generic approaches (Palm et al., 2019). There are some efforts in other domains such as dental image analysis (Krois et al., 2021). Research has also been conducted in the context of DLA, where the authors advocate the combination of novel models and curated (synthetic) datasets to address the challenge of generalizability (Naiman, 2023). However, there are not many efforts in the context of KIE.
  6. (6)
    Domain knowledge: As discussed in chapter 6.2, only a few authors integrate domain knowledge into the proposed approaches and/or adopt a practice-oriented view in general. However, it seems promising to explicitly integrate existing domain knowledge into KIE systems, which is also shown in (Cui et al., 2021), as it contains valuable information about the corresponding business processes, their documents and how they need to be understood as a whole. The design of novel pre-training tasks could be one possibility for integrating these aspects.
  7. (7)
    Real-world usability: The relevance of automated document processing in real-world scenarios is indisputable (see also the review by (Martínez-Rojas et al., 2023)). Because of this great importance, more efforts should be made to improve the overall practicability of KIE approaches. The previously mentioned research directions already address this matter. For example, one could consider pre-processing techniques to remove document image artifacts that are specific to real-world settings, such as removing stamps (Yang et al., 2023b). The aforementioned focus on generalization capabilities can also have a positive impact on the practicality of KIE systems since deploying a system that can properly process many different types of documents simultaneously could result in lower costs and less maintenance.
    On top of that, additional aspects could be considered. For example, (Kivimäki et al., 2023) propose a method to provide confidence estimates for the extracted data. The authors emphasise that in industrial settings, the primary goal is typically to make decisions based on model predictions rather than the raw extracted data. Research in this direction could greatly assist the collaboration between KIE systems and human administrators during document processing tasks, as it is essential to validate automatically extracted data in real-world settings (Houy et al., 2019). (Sassioui et al., 2023) also emphasize the importance of real-time KIE in industry applications. In this regard, a focus on lightweight KIE solutions should be considered. This could for example include investigating the trade-off between model size and extraction performance more closely. (Hamri et al., 2023) have shown that a relatively small model with around 50 thousand parameters can produce competitive results compared to systems that consist of several hundred million parameters.
  8. (8)
    End-to-end performance: The evaluation has shown that end-to-end approaches usually achieve inferior extraction results compared to systems that use an external OCR engine. In this regard, more research should be conducted on how to improve the text recognition step, as the end-to-end systems have a high potential due to their independence of OCR engines, also highlighted by (Sassioui et al., 2023). End-to-end approaches have received less attention so far, which is also reflected in the number of identified approaches. Nevertheless, there is a relatively strong increase in corresponding methods, especially since 2022.
8. Conclusion
Research in the area of Key Information Extraction has seen an increased interest in recent years, mainly due to major advances in the field of Deep Learning. Nowadays, even visually-rich business documents with very complex layouts and high information-richness can be automatically processed by corresponding systems. To this end, this manuscript represents a systematic literature review covering the research on Key Information Extraction between 2017 and 2023 including 96 proposed approaches, with the aim of identifying the state of the art in this field and identifying potentials for further research. The identified methods were compared both qualitatively and quantitatively based on various characteristics.
The analysis has shown that related work tends to follow a very narrow corridor in which already proven concepts are being successively refined and improved. In general, the approaches follow three key paradigms for representing document images, namely as sequences, graphs or grids. In detail, the approaches differ in their choice of architectures used for encoding and decoding the input documents, although it can also be seen that certain models are used more frequently (e.g., BERT-based models for textual inputs). Besides, novel concepts have been explored over time, such as OCR-independent and autoregressive methods, which on the one hand do not require an external OCR engine for preliminary text reading steps, and on the other hand can output arbitrary text, making them more flexible in terms of the downstream tasks they support. The authors investigate how different input modalities implicitly and explicitly contained in complex documents can be optimally integrated into the model architectures. In particular, visual cues that can be obtained from document images are increasingly incorporated into the models in order to improve their performance for more complex use cases. Much effort is also put into learning a general document understanding through deep learning models, which can then be used for Key Information Extraction. This is usually done by extensive and innovative pre-training tasks followed by specific fine-tuning steps. Another general observation is that the complexity of the corresponding models in terms of the number of parameters does not play a significant role and that even lightweight models can achieve promising extraction results.
The research area is strongly characterized by the fact that an improvement of the state of the art is achieved by obtaining better extraction results on established benchmark datasets. However, a quantitative comparison of the presented results is not always meaningful, since the evaluation setups used by the authors are very heterogeneous. Furthermore, for most of the benchmarks, very good results have already been achieved (F1-scores above 0.97). The research area should therefore move away from focusing on improving benchmarks results and instead investigate innovations along other dimensions. Future work could focus on even more lightweight models in order to improve their practical applicability. It could also be investigated how the models can be designed in order to require less data for training procedures. We identified several other starting points for follow-up research based on the findings. These include proposing novel and more diverse datasets as well as consistent evaluation setups that allow for an adequate quantitative comparison. We also advocate that more attention should be paid to the real-world usability of corresponding approaches and to the integration of domain knowledge, since on the one hand document processing tasks play a key role in daily business workloads and, on the other hand, offer a special perspective on Key Information Extraction with unique requirements, but also possibilities.
A. Business process perspective
Figure SF1. Exemplary run of a purchasing process
[…]
A graph that shows a run of a purchasing process with document-based exchanges highlighted. Refer to caption
B. Database search strings
The title should include the name of the DU subtask, i.e., KIE, its variant formulations, and other commonly used terms such as entity extraction. We also included document* in the title search. This in particular resulted in many non-relevant papers, but is important, as there are many other terms used for this research area, as mentioned before. This increased the manual effort for filtering, but also increased the recall and thus the coverage. Furthermore, titles, abstracts, keywords or metadata should include terms such as visually rich or image in order to find literature that specifically deals with VRDs according to the scope of this work. We also include terms such as invoice or receipts, since these are the most common terms regarding document types. However, these terms were included with OR-operators, which means that the results do not necessarily have to include them. To identify approaches using DL methods, we added the terms deep learning, artificial intelligence and neural network*, where at least one of them should appear somewhere in the text or metadata. In order to filter out papers that are outside the scope of this work, we included terms that should not appear in the title, namely handwrit*, histor*, table or web, if supported by the search engine. We also used wildcards (*) to allow for different expressions of the search terms (e.g., (information AND extract*)).
Table ST1. Defined search strings
DB Search string Hits
ACL intitle:((information AND extract*) OR ”entity extraction” OR ”entity recognition” OR document* OR ”visually rich”) AND (intitle:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*) OR abstract:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*) OR keywords:(”visually?rich” OR modal* OR image* OR invoice* OR receipt*)) 90
ACM Title:((information AND extract*) OR ”entity extraction” OR ”entity recognition” OR document* OR ”visually?rich”) AND ( Title:(”visually?rich” OR modal* OR image OR invoice OR receipt) OR Abstract:(”visually?rich” OR modal* OR image OR invoice OR receipt) OR Keyword:(”visually?rich” OR modal* OR image OR invoice OR receipt) ) AND NOT (Title:(handwrit*) OR Title:(histor*) OR Title:(table) OR Title:(web)) 248
AIS ((title:information AND title:extract*) OR title:”entity extraction” OR title:”entity recognition” OR title:document* OR title:”visually rich”) AND (”deep learning” OR ”artificial intelligence” OR ”neural network*”) &start_date=01/01/2017&end_date=12/31/2023 28
IEEE (((”Document Title”:information AND ”Document Title”:extract*) OR ”Document Title”:”entity extraction” OR ”Document Title”:”entity recognition” OR ”Document Title”:document* OR ”Document Title”:”visually rich”) AND ( (”Document Title”:”visually rich” OR ”Document Title”:modal* OR ”Document Title”:invoice OR ”Document Title”:receipt) OR (”Abstract”:”visually rich” OR ”Abstract”:modal* OR ”Abstract”:invoice OR ”Abstract”:receipt) OR (”Author Keywords”:”visually rich” OR ”Author Keywords”:modal* OR ”Author Keywords”:invoice OR ”Author Keywords”:receipt) ) NOT (”Document Title”:handwrit* OR ”Document Title”:histor* OR ”Document Title”:table)) 98
SD Year: 2017-2023; Title, abstract, keywords: ”visually rich” OR modal OR image OR invoice OR receipt; Title: (information AND extract) OR document OR ”visually rich” ; ”deep learning” OR ”artificial intelligence” OR cognitive OR intelligent OR ”neural network” 163
SL query=”visually+rich”+OR+modal*+OR+image+OR+invoice+OR+receipt & dc.title=information+extract* & date-facet-mode=between&facet-start-year=2017 & facet-end-year=2023 269
Total 896
C. Numbering of the approaches
The manuscripts are sorted chronologically by year of publication. Within a year, they are sorted alphabetically by author name.
Table ST2. Numbering of the approaches
# Reference Title
1 (Palm et al., 2017) (2017) CloudScan - A Configuration-Free Invoice Analysis System Using Recurrent Neural Networks
2 (Katti et al., 2018) (2018) Chargrid: Towards understanding 2D documents
3 (Denk and Reisswig, 2019) (2019) BERTgrid: Contextualized Embedding for 2D Document Representation and Understanding
4 (Gal et al., 2019) (2019) Visual-Linguistic Methods for Receipt Field Recognition
5 (Guo et al., 2019) (2019) EATEN: Entity-aware attention for single shot visual text extraction
6 (Liu et al., 2019a) (2019a) Graph convolution for multimodal information extraction from visually rich documents
7 (Lohani et al., 2019) (2019) An Invoice Reading System Using a Graph Convolutional Network
8 (Palm et al., 2019) (2019) Attend, copy, parse end-to-end information extraction from documents
9 (Qian et al., 2019) (2019) GraphIE: A Graph-Based Framework for Information Extraction
10 (Carbonell et al., 2020) (2020) Named entity recognition and relation extraction with graph neural networks in semi structured documents
11 (Cheng et al., 2020) (2020) One-shot Text Field labeling using Attention and Belief Propagation for Structure Information Extraction
12 (Dang and Thanh, 2020) (2020) End-to-end information extraction by character-level embedding and multi-stage attentional u-net
13 (Gal et al., 2020) (2020) Cardinal Graph Convolution Framework for Document Information Extraction
14 (Hua et al., 2020) (2020) Attention-Based Graph Neural Network with Global Context Awareness for Document Understanding
15 (Luo et al., 2020) (2020) Merge and Recognize: A Geometry and 2D Context Aware Graph Model for Named Entity Recognition from Visual Documents
16 (Majumder et al., 2020) (2020) Representation learning for information extraction from form-like documents
17 (Oral et al., 2020) (2020) Information Extraction from Text Intensive and Visually Rich Banking Documents
18 (Sage et al., 2020) (2020) End-to-End Extraction of Structured Information from Business Documents with Pointer-Generator Networks
19 (Wang et al., 2020) (2020) DocStruct: A multimodal method to extract hierarchy structure in document for general form understanding
20 (Wei et al., 2020) (2020) Robust Layout-aware IE for Visually Rich Documents with Pre-trained Language Models
21 (Xu et al., 2020) (2020) LayoutLM: Pre-training of Text and Layout for Document Image Understanding
22 (Yu et al., 2020) (2020) PICK: Processing Key Information Extraction from Documents using Improved Graph Learning-Convolutional Networks
23 (Zhang et al., 2020) (2020) TRIE: End-to-End Text Reading and Information Extraction for Document Understanding
24 (Appalaraju et al., 2021) (2021) DocFormer: End-to-End Transformer for Document Understanding
25 (Belhadj et al., 2021) (2021) Consideration of the Word’s Neighborhood in GATs for Information Extraction in Semi-structured Documents
26 (Davis et al., 2021) (2021) Visual FUDGE: Form Understanding via Dynamic Graph Editing
27 (Garncarek et al., 2021) (2021) LAMBERT: Layout-Aware Language Modeling for Information Extraction
28 (Gu et al., 2021) (2021) Unified Pretraining Framework for Document Understanding
29 (Hamdi et al., 2021) (2021) Information Extraction from Invoices
30 (Holeček, 2021) (2021) Learning from similarity and information extraction from structured documents
31 (Hwang et al., 2021a) (2021a) Cost-effective End-to-end Information Extraction for Semi-structured Document Images
32 (Hwang et al., 2021b) (2021b) Spatial Dependency Parsing for Semi-Structured Document Information Extraction
33 (Kerroumi et al., 2021) (2021) VisualWordGrid: Information Extraction from Scanned Documents Using a Multimodal Approach
34 (Klaiman and Lehne, 2021) (2021) DocReader: Bounding-Box Free Training of a Document Information Extraction Model
35 (Krieger et al., 2021) (2021) Information Extraction from Invoices: A Graph Neural Network Approach for Datasets with High Layout Variety
36 (Lee et al., 2021) (2021) ROPE: Reading Order Equivariant Positional Encoding for Graph-based Document Information Extraction
37 (Li et al., 2021a) (2021a) StructuralLM: Structural pre-training for form understanding
38 (Li et al., 2021b) (2021b) SelfDoc: Self-Supervised Document Representation Learning
39 (Li et al., 2021c) (2021c) StrucTexT: Structured Text Understanding with Multi-Modal Transformers
40 (Lin et al., 2021) (2021) ViBERTgrid: A Jointly Trained Multi-modal 2D Document Representation for Key Information Extraction from Documents
41 (Nguyen et al., 2021) (2021) A Span Extraction Approach for Information Extraction on Visually-Rich Documents
42 (Ning et al., 2021) (2021) A Segment-Based Layout Aware Model for Information Extraction on Document Images
43 (Powalski et al., 2021) (2021) Going Full-TILT Boogie on Document Understanding with Text-Image-Layout Transformer
44 (Sarkhel and Nandi, 2021) (2021) Improving information extraction from visually rich documents using visual span representations
45 (Tang et al., 2021) (2021) MatchVIE: Exploiting Match Relevancy between Entities for Visual Information Extraction
46 (Tata et al., 2021) (2021) Glean: Structured extractions from templatic documents
47 (Wang et al., 2021a) (2021a) Towards Robust Visual Information Extraction in Real World: New Dataset and Novel Solution
48 (Wang et al., 2021b) (2021b) Tag, Copy or Predict: A Unified Weakly-Supervised Learning Framework for Visual Information Extraction using Sequences
49 (Xu et al., 2021) (2021) LayoutLMv2: Multi-modal pre-training for visually-rich document understanding
50 (Zhang et al., 2021) (2021) Entity Relation Extraction as Dependency Parsing in Visually Rich Documents
51 (Arroyo et al., 2022) (2022) Key Information Extraction in Purchase Documents using Deep Learning and Rule-based Corrections
52 (Cao et al., 2022a) (2022a) Query-driven Generative Network for Document Information Extraction in the Wild
53 (Cao et al., 2022b) (2022b) GMN: Generative Multi-modal Network for Practical Document Information Extraction
54 (De Trogoff et al., 2022) (2022) Automatic Key Information Extraction from Visually Rich Documents
55 (Du et al., 2022) (2022) CALM: Commen-Sense Knowledge Augmentation for Document Image Understanding
56 (Gu et al., 2022) (2022) XYLayoutLM: Towards Layout-Aware Multimodal Networks For Visually-Rich Document Understanding
57 (Hong et al., 2022) (2022) BROS: A Pre-trained Language Model Focusing on Text and Layout for Better Key Information Extraction from Documents
58 (Huang et al., 2022) (2022) LayoutLMv3: Pre-training for Document AI with Unified Text and Image Masking
59 (Kim et al., 2022) (2022) OCR-Free Document Understanding Transformer
60 (Lee et al., 2022) (2022) FormNet: Structural Encoding beyond Sequential Modeling in Form Document Information Extraction
61 (Li et al., 2022c) (2022c) Relational Representation Learning in Visually-Rich Documents
62 (Oral and Eryiğit, 2022) (2022) Fusion of visual representations for multimodal information extraction from unstructured transactional documents
63 (Peng et al., 2022) (2022) ERNIE-Layout: Layout Knowledge Enhanced Pre-training for Visually-rich Document Understanding
64 (Wang et al., 2022b) (2022b) LiLT: A Simple yet Effective Language-Independent Layout Transformer for Structured Document Understanding
65 (Wang et al., 2022a) (2022a) MmLayout: Multi-grained MultiModal Transformer for Document Understanding
66 (Yeghiazaryan et al., 2022) (2022) Tokengrid: Toward More Efficient Data Extraction From Unstructured Documents
67 (Zhang et al., 2022a) (2022a) Extract Data Points from Invoices with Multi-layer Graph Attention Network and Named Entity Recognition
68 (Zhang et al., 2022b) (2022b) Dual-VIE: Dual-Level Graph Attention Network for Visual Information Extraction
69 (Belhadj et al., 2023a) (2023a) Improving Information Extraction from Semi-structured Documents Using Attention Based Semi-variational Graph Auto-Encoder
70 (Belhadj et al., 2023b) (2023b) Low-Dimensionality Information Extraction Model for Semi-structured Documents
71 (Cao et al., 2023) (2023) GenKIE: Robust Generative Multimodal Document Key Information Extraction
72 (Davis et al., 2023) (2023) End-to-End Document Recognition and Understanding with Dessurt
73 (Deng et al., 2023b) (2023b) An Iterative Graph Learning Convolution Network for Key Information Extraction Based on the Document Inductive Bias
74 (Deng et al., 2023a) (2023a) GenTC: Generative Transformer via Contrastive Learning for Receipt Information Extraction
75 (Dhouib et al., 2023) (2023) DocParser: End-to-end OCR-Free Information Extraction from Visually Rich Documents
76 (Do et al., 2023) (2023) A Novel Approach for Extracting Key Information from Vietnamese Prescription Images
77 (Gbada et al., 2023) (2023) VisualIE: Receipt-Based Information Extraction with a Novel Visual and Textual Approach
78 (Gemelli et al., 2023) (2023) Doc2Graph: A Task Agnostic Document Understanding Framework Based on Graph Neural Networks
79 (Guo et al., 2024) (2024) DCMAI: A Dynamical Cross-Modal Alignment Interaction Framework for Document Key Information Extraction
80 (Hamri et al., 2023) (2023) Enhancing GNN Feature Modeling for Document Information Extraction Using Transformers
81 (He et al., 2023b) (2023b) ICL-D3IE: In-Context Learning with Diverse Demonstrations Updating for Document Information Extraction
82 (He et al., 2023c) (2023c) Document Information Extraction via Global Tagging
83 (Kuang et al., 2023) (2023) Visual Information Extraction in the Wild: Practical Dataset and End-to-End Solution
84 (Lee et al., 2023) (2023) FormNetV2: Multimodal Graph Contrastive Learning for Form Document Information Extraction
85 (Li et al., 2023b) (2023b) Enhancing Visually-Rich Document Understanding via Layout Structure Modeling
86 (Liao et al., 2023) (2023) DocTr: Document Transformer for Structured Information Extraction in Documents
87 (Luo et al., 2023) (2023) GeoLayoutLM: Geometric Pre-training for Visual Information Extraction
88 (Shi et al., 2023) (2023) Multi-scale Cell-based Layout Representation for Document Understanding
89 (Tang et al., 2023) (2023) Unifying Vision, Text, and Layout for Universal Document Processing
90 (Tu et al., 2023) (2023) LayoutMask: Enhance Text-Layout Interaction in Multi-modal Pre-training for Document Understanding
91 (Wang et al., 2023a) (2023a) DocGraphLM: Documental Graph Language Model for Information Extraction
92 (Yang et al., 2023a) (2023a) Modeling Entities as Semantic Points for Visual Information Extraction in the Wild
93 (Yu et al., 2023) (2023) StrucTexTv2: Masked Visual-Textual Prediction for Document Image Pre-training
94 (Zhang et al., 2023b) (2023b) Reading Order Matters: Information Extraction from Visually-rich Documents by Token Path Prediction
95 (Zhang et al., 2023a) (2023a) A Character-Level Document Key Information Extraction Method with Contrastive Learning
96 (Zhang et al., 2023c) (2023c) Multimodal Pre-training Based on Graph Attention Network for Document Understanding
D. Overview
The individual papers were examined according to the following general properties, each of which considers different aspects, also with practical applications in mind.
  • Category: Indicates the superordinate method (see section 2.3), or a combination thereof, used specifically for the KIE task, if applicable.
  • Input modalities: Indicates, which modalities derived from input documents were considered for KIE. Besides textual, visual and layout-oriented features, we also include custom hand-crafted features in the analysis.
  • Data basis: We gather the amount as well as types of documents used to implement and evaluate the proposed approaches, which provides an assessment of the underlying data basis and data requirements.
  • Reproducibility & deployability: These aspects highlight the availability of the implemented artifacts in terms of implemented code and/or model weights and whether they allow to be used in commercial settings.
  • Independent of external OCR: Indicates whether the KIE approach requires no external OCR engine. If this is the case, the system is either responsible both for text reading and information extraction (end-to-end) or alternatively requires no text reading stage at all and maps input documents immediately to desired outputs (OCR-free) (Sassioui et al., 2023).
  • Number of parameters: If specified by the authors, indicates the total number of model parameters (in millions) and thus the overall complexity of the model. If multiple model variants are proposed, the largest ones are listed.
  • Integration of domain knowledge: Indicates whether the authors integrate domain knowledge (e.g., knowledge about business processes or values to be extracted) in some form.
  • Evolution of existing KIE approach: Indicates whether the proposed approach is a successor to an already existing KIE method with distinct refinements.
  • Evaluation in industry setting: Indicates whether the proposed approach has been evaluated in real-world industry settings - either in a quantitatively or qualitatively manner. This does not include cases where authors simply report the runtime of the approaches.
  • Weakly annotated: Makes statements about the annotations required for the approach, i.e., whether fully-annotated documents including word-level annotations and bounding boxes are required or not. This therefore gives an indication of the effort required to train the corresponding approaches.
  • Generative: Indicates whether the KIE system is an autoregressive approach that can produce arbitrary output during decoding steps.
Table ST3. Overview of the analysis
# Category Input modalities Data basis Reproducibility & deployability Independent of external OCR Number of parameters (in M) Integration of domain knowledge Evolution of existing KIE approach Evaluation in industry setting Weakly annotated Generative
text
image
layout
hand-crafted
number of documents used for KIE implementation
Document types for KIE implementation
Code available
Model weights available
Commercial use
1 sequence 326471 Invoices n/a
[…]
4 grid 5094 Receipts n/a
[…]
6 graph 4500 Invoices, receipts n/a
7 graph 3100 Invoices n/a
8 grid 1181990 Invoices n/a
9 graph 25200 Reports n/a
10 graph 324 Forms, historical documents 201
11 graph 1407 Receipts, certificates, administrative documents n/a
12 grid 461 Invoices, receipts 1,05
13 graph 7485 Invoices, receipts n/a
14 graph 2089 Contracts, receipts n/a
15 graph 14471 Receipts, business cards n/a
[…]
19 1056 Medical reports, forms n/a
20 graph + sequence 2472 Invoices, resumes n/a
[…]
22 graph + sequence 5533 Invoices, tickets, receipts n/a
23 8448 Invoices, receipts, resumes n/a
24 sequence 1540 Receipts, NDAs $\sim similar-to\sim$ 536
25 graph 3453 Invoices, payslips, receipts n/a
26 graph 199 Forms 17
27 sequence 4944 NDAs, reports, receipts 125
28 sequence 1199 Forms, receipts 274
29 sequence 154047 Invoices n/a
30 graph + sequence 25701 Invoices n/a
31 sequence 133793 Name cards, receipts n/a
32 graph 37605 Forms, receipts, invoices, namecards n/a
[…]
35 graph 1129 Invoices n/a
36 graph 18199 Invoices, forms n/a
37 sequence 199 Forms 355
38 sequence 199 Forms 146
39 sequence 2666 Forms, receipts, exams 107
[…]
41 sequence 2200 Invoices, receipts 113
42 sequence 2712 Forms, receipts n/a
[…]
45 graph + sequence 2666 Forms, exams, receipts n/a
[…]
47 sequence 2467 Reciepts, exams n/a
48 grid + sequence 4798 Receipts, exams, business licenses n/a
49 sequence 2712 Forms, receipts, NDAs 426
50 graph + sequence 1799 Forms, customs declarations n/a
[…]
52 sequence 7865 Forms, receipts, exams, miscellaneous n/a
53 sequence 2172 Forms, receipts n/a
54 graph 973 Receipts n/a
55 graph + sequence 199 Forms 285
56 sequence 1592 Forms n/a
57 sequence 1825 Forms, receipts 340
58 sequence 1199 Forms, receipts 368
59 sequence 65500 Receipts, tickets, business cards 143
60 graph + sequence 11199 Forms, receipts, invoices 217
61 sequence 1199 Forms, receipts 142
[…]
63 sequence 2513 Forms, receipts, NDAs n/a
64 sequence 2693 Forms, receipts, exams n/a
65 graph + sequence 2172 Forms, receipts 426
66 grid 8700 Invoices 20,5
67 graph 513 Invoices n/a
68 graph 4699 Forms, bills n/a
[…]
70 graph 3973 Invoices, receipts 41
71 sequence 2172 Forms, receipts 180
72 sequence 1867 Forms, handwritten documents 127
73 graph 2740 Receipts 66
74 sequence 32824 Receipts, exams, invoices n/a
75 sequence 15973 Receipts, information statements 70
76 grid 10693 Prescriptions n/a
77 sequence 1740 Receipts 140
78 graph 199 Forms 6,2
79 sequence 1739 Forms, receipts, NDAs 362
80 graph 973 Receipts 0,0536
81 sequence 2712 Forms, Receipts 175000
82 sequence 1592 Forms n/a
83 sequence 3973 Receipts n/a
84 graph + sequence 17004 Forms, receipts, NDAs 204
85 graph + sequence 1398 Forms, receipts 372
86 sequence 1199 Forms, receipts 153
87 sequence 1199 Forms, receipts 399
88 1199 Forms, receipts 368
89 sequence 1199 Forms, receipts 794
90 sequence 2712 Forms, receipts 404
91 graph 1199 Forms, receipts n/a
92 sequence 2398 Invoices, bills, receipts 50
93 sequence 199 Forms 238
94 graph 1199 Forms, receipts n/a
95 sequence 1172 Receipts 35
96 graph 2172 Forms, receipts 265
E. Sequence-based approaches
Table ST4. Analysis of sequence-based approaches
# Granularity Pre-Training Textual encoder Visual encoder Decoder Special attention mechanism
# documents used for Pre-Training Datasets Tasks
1 word / / / / / LSTM
5 character / / / / Inceptionv3 LSTM Entity-aware attention mechanism
17 word + character / / / FastText, ELMo, BERT / BiLSTM-CRF
18 word / / / BiLSTM / LSTM Global attention
20 word / / / RoBERTa / Sequence labeling layer
21 word 6000000 IIT-CDIP Masked Visual-Language Modeling, Multi-label Document Classification BERT Faster R-CNN (ResNet101) Sequence labeling layer Self-attention with absolute 2-D position features
22 character / / / Transformer ResNet50 BiLSTM-CRF /
24 word 5000000 IIT-CDIP* Multi-Modal Masked Language Modeling, Learn To Reconstruct, Text Describes Image LayoutLM ResNet50 Sequence labeling layer Multi-modal self-attention
27 word 315000 Common Crawl* Masked language modeling RoBERTa / Sequence labeling layer Gated Cross-Attention
28 sentence 1000000 IIT-CDIP* Masked Sentence Modeling, Visual Contrastive Learning, Vision-Language Alignment SBERT ResNet50 Sequence labeling layer Gated Cross-Attention
29 word / / / CamemBERT / CRF
30 word / / / / CNN Sequence labeling layer
31 word 6000000 Custom n/a BERT / Transformer + gated copying mechanism + parser
34 word / / / / Chargrid-OCR LSTM
37 cell 6000000 IIT-CDIP Masked Visual-Language Modeling, Cell Position Classification BERT / Sequence labeling layer
38 sentence 320000 RVL-CDIP Masked Visual-Language Modeling SBERT Faster R-CNN Sequence labeling layer Modality-adaptive attention
39 sentence 900000 DocBank, RVL-CDIP Masked Visual-Language Modeling, Segment Length Prediction, Paired Box Direction ERNIE ResNet50 Sequence labeling layer
41 word 5170 Custom Span extraction semantic pre-training LayoutLM / Sequence labeling layer
42 segment / / / BERT / BiLSTM-CRF
43 word 1105000 RVL-CDIP*, UCSF Industry Documents Library, Common Crawl* Masked Visual-Language Modeling T5 U-Net n/a With relative attention biases
45 segment / / / BERT ResNet34 MLP + CRF
47 segment + token / / / n/a Mask R-CNN (ResNet50) BiLSTM-CRF two multi-layer multi-head self-attention modules
48 word / / / Custom ResNet18, U-Net BiLSTM
49 word 6000000 IIT-CDIP Masked Visual-Language Modeling, Text-Image Alignment, Text-Image Matching UniLMv2 ResNeXt101-FPN Sequence labeling layer Spatial-Aware Self-Attention
50 word / / / word2vec, BERT, LayoutLM, BiLSTM / MLP + BiLSTM
51 word / / / Transformer CNN n/a
52 word 43000000 IIT-CDIP*, Custom Unidirectional LM, Bidirectional LM, Sequence-to-Sequence LM LayoutXLM n/a Custom
53 word 6000000 IIT-CDIP Unidirectional LM, Bidirectional LM, Sequence-to-Sequence LM, NER-LM UniLM, MD-BERT ResNet18 MD-BERT Modal-aware multi-head attention
55 word / / / LayoutLMv2 LayoutLMv2 MLP
56 word / / / LayoutXLM ResNeXt101-FPN Bi-affine classifier
57 word 6000000 IIT-CDIP Token-masked LM, area-masked LM BERT / Sequence labeling layer, SPADE
58 word 6000000 IIT-CDIP Masked Language Modeling, Masked Image Modeling, Word-Patch Alignment RoBERTa DiT MLP
59 / 8000000 IIT-CDIP, Custom ”pseudo-OCR” / Swin Transformer BART
60 word 700000 Custom Masked Language Modeling ETC / Viterbi algorithm Rich Attention (also considers spatial distance)
61 word 900000 DocBank, RVL-CDIP Masked Visual-Language Modeling, Relational Consistency Modeling LayoutLMv2 ResNet50-FPN Custom
62 word / / / ELMo CNN, LGCN n/a
63 word 10000000 IIT-CDIP* Reading Order Prediction, Replaced Region Prediction, Masked Visual-Language Modeling, Text-Image Alignment RoBERTa Faster-RCNN Sequence labeling layer Spatial-aware Disentangled Attention
64 word 6000000 IIT-CDIP Masked Visual-Language Modeling, Key Point Location, Cross-modal Alignment Identification RoBERTa, InfoXLM / Sequence labeling layer Bidirectional attention complementation mechanism
65 word + segment / / / LayoutLMv2 LayoutLMv2 Sequence labeling layer Spatial-Aware Self-Attention
71 word / / / OFA OFA OFA + parser
72 / ¿ 6000000 IIT-CDIP, Custom Over 25 different tasks / Swin Transformer Greedy search decoding Cross-attention to entire to entire visual array
74 word / / / LayoutXLM Donut Donut & Greedy search decoding
75 / 1700000 IIT-CDIP* Knowledge Transfer, Masked Document Reading / ConvNeXt, Swin Transformer Transformer
77 word / / / N-gram-model VGG16 Sequence labeling layer
79 word 400000 RVL-CDIP Joint representation learning, crossover alignment, masked text reconstruction loss BERT ViT n/a
81 word / / / / / Davinci-003 & ChatGPT
82 word / / / LayoutXLM LayoutXLM Learned tagging matrix and custom decoding
83 text line / / / / Mask R-CNN (ResNet50-FPN) LSTM-CRF Self-Attention in custom encoder
84 word 6000000 IIT-CDIP Masked Language Modeling, Graph Contrastive Learning ETC ConvNet ETC
85 word 6000000 IIT-CDIP Masked Language Modeling, Masked Image Modeling, Word-Patch Alignment LayoutLMv3 LayoutLMv3 Sequence labeling layer Spatial-Aware Self-Attention & graph-aware Transformer layer
86 word 6000000 IIT-CDIP Masked Detection Modeling BERT ResNet-50, Deformable DETR Custom Vision-Language Decoder Deformable cross-attention & language-conditioned cross-attention
87 word 6000000 IIT-CDIP Masked Visual-Language Modeling, Direction and Distance Modeling, Detection of Direction Exceptions, Collinearity Identification of Triplet BROS ConvNeXt-FPN Coarse Relation Prediction & Relation Feature Enhancement
89 word ¿ 6000000 IIT-CDIP, DocBank, Kleister Charity, PWC, DeepForm Layout Modeling, Visual Text Recognition, Joint Text-Layout Reconstruction, Masked Image Reconstruction, Classification, Layout Analysis, Information Extraction, Question Answering, Document NLI Transformer Transformer Text-Layout Decoder & MAE
90 segment 10000000 IIT-CDIP* Whole Word Masking, Layout-AwareMasking, Masked Position Modeling XLM-RoBERTa / n/a
92 / 900000 DocBank, RVL-CDIP Entity-Image Text Matching / ConvNeXt-FPN Sequence labeling layer
93 / 11000000 IIT-CDIP Masked Language Modeling, Masked Image Modeling (with text region-level masking) / ResNet50, ResNeXt-101 Differentiable Binarization + Transformer + MLP
95 text line / / / CharLM, Custom / BiLSTM-CRF
F. Graph-based approaches
Table ST5. Analysis of graph-based approaches
# Graph creation Nodes Edges Graph propagation Method of KIE
Granularity Features Feature vector size Global node Direction Features Propagation type Layers
6 Fully connected segment BiLSTM-embeddings n/a directed Horizontal and vertical distance + aspect ratio + relative height + relative width GCN on node-edge-node-triplets with MLP and self-attention 2 BiLSTM-CRF
7 Nearest neighbors in 4 major directions word Boolean features + relative distance to nearest neighbors + Byte Pair Encoding 317 undirected / GCN with fourier transformation 5 Softmax layer (node classification)
9 Nearest neighbors in 4 major directions sentence BiLSTM-embeddings + CharCNN-embeddings 96 directed or undirected Type (left-to-right, right-to-left, up-to-down, down-to-up) GCN 1 BiLSTM-CRF
10 k-NN (k=10) & fully connected word Fasttext-embeddings + normalized coordinates n/a undirected / GAT n/a MLP
11 Landmark and field connected + pairwise connection between fields (36 rays emitted) word / / directed Normalized coordinates MLP + CRF / Belief Propagation
13 Nearest neighbors in 4 major directions word 1-hot encoded adjacency tensor + miscellaneous n/a undirected / Cardinal GCN with mapping matrix n/a Softmax layer (node classification)
14 Segments in same and previous line connected segment BERT-embeddings + ResNet50-embeddings 1024 directed / GAT + multi-head attention 3 CRF
15 Fully connected word BiLSTM-embeddings + relative coordinates + relative text size 304 undirected / GCN + custom merge layer 1 LSTM-CRF
20 Closest horizontal or vertical neigbor sentence RoBERTa-embeddings + font-encoding 776 undirected / GCN 2 Sequence labeling layer
22 Learned soft adjacency matrix sentence Transformer-embeddings + ResNet50-embeddings 1024 undirected Horizontal and vertical distance + aspect ratio + relative height + relative width + sentence length ratio GCN on node-edge-node-triplets with MLP 1 BiLSTM-CRF
25 Groups all words within limited horizontal and vertical distance word Boolean features + normalized coordinates + Byte Pair Encoding 310 undirected / GAT + multi-head attention 3 Softmax layer (node classification)
26 Iteratively learned via NN text lines CNN-encodings + normalized coordinates + detection confidence + class prediction 256 directed CNN-embeddings + normalized coordinates + class prediction + distance GCN + multi-head attention 7;7;4 Node classification
30 Closest neigbor in each 90 degree angle (total of 4 neigbors for each word) word Statistical text features + 1-hot encoded characters + image-crop 640 directed / GCN n/a Custom
32 Learned binary matrix word BERT-embeddings + relative coordinates + physical distance + relative angle n/a directed Type (rel-s, rel-g) / / Custom (seeding, serialization, grouping)
35 Nearest neighbors in 4 major directions + self-loop word 1-hot encoded characters + normalized coordinates + BERT-embeddings 110x10 & 768 & 13 undirected / GAT 3 Softmax layer (node classification)
36 ß-skeleton graph (ß=1) (custom reading order equivariant ) word BERT-embeddings + normalized coordinates n/a directed Normalized relative distances + MobileNetV3-embeddings MLP + GCN 5 Node classification
45 Fully connected text lines BERT-embeddings + custom Num2Vec + ResNet34-embeddings 768 n/a Num2Vec-embedding of: horizontal and vertical distance + aspect ratio + relative height + relative width GCN on node-edge-node-triplets with MLP and self-attention 2 MLP+CRF + BiLSTM
50 Fully connected word Word-embeddings (Word2Vec/BERT/LayoutLM) + label-embedding 200 or 864 n/a Horizontal and vertical distance GCN on node-edge-node-triplets with MLP and self-attention 2 MLP + biaffine attention
54 k-NN (k=5) segment Average pooling of Fasttext-embeddings + boolean features + ResNet34-embeddings n/a undirected / GAT 3 BiLSTM-CRF
60 ß-skeleton graph (ß=1) word One-hot BERT-embeddings + normalized coordinates n/a directed Relative distance + shortest distance + aspect ratio GCN 12 Viterbi algorithm
65 Fully connected & custom (between fine- and coarse-grained) word + segment LayoutLMv2-embeddings + common-sense-embedding n/a undirected Type (fine-grained, cross-grained, coarse-grained) GCN with spatial-aware self-attention n/a Sequence labeling layer
67 Fully connected segment Byte Pair Encoding + normalized coordinates + label-embedding (NER category) 128 n/a / GAT + multi-head attention 2 MLP
68 Fully connected subgraph for each bounding box; k-NN (k=4) token + text lines LayoutLM-embeddings + normalized coordinates + SBERT-embeddings + ResNet101-embeddings n/a undirected / GAT + circuit-breaking attention 2 BiLSTM-CRF
69 k-NN in three lines (same line, line above and line below if existing); k = 4&8; max 256 nodes segment Byte Pair Encoding + normalized coordinates + region encoding + ResNet-50+U-Net-embeddings 337 undirected / GAT + multi-head attention 4&6 Softmax layer (node classification) + VGAE
70 k-NN in three lines (same line, line above and line below if existing); max 256 nodes word Byte Pair Encoding + normalized coordinates + region encoding + ResNet-50+UNet-embeddings n/a undirected / GAT + multi-head attention 4 Softmax layer (node classification)
73 k-NN (k=15) segment Word2Vec-embeddings + ResNet50-embeddings 512 undirected Relative coordinates GCN on node-edge-node-triplets with MLP and self-attention 2 BiLSTM-CRF
78 Fully connected and iteratively learned via NN word or segment spaCy-embeddings + U-Net-embeddings + normalized coordinates n/a directed or undirected Normalized distance + one-hot encoded relative coordinates Modified GraphSAGE 1 Node and edge predictor (FC)
80 k-NN in three lines (same line, line above and line below if existing); k = 4 word LayoutLMv2-embeddings + normalized coordinates + custom features 777 undirected / Graph transformer 4 Softmax layer (node classification)
84 ß-skeleton graph (ß=1) word One-hot BERT-embeddings n/a directed Relative distance + shortest distance + aspect ratio + ConvNet-embeddings GCN 6 ETC
85 Custom layout tree; nearest neighbors in 4 major directions text lines / n/a directed / Custom graph reordering & graph masking 1 Sequence labeling layer
91 Closest neigbor in each 45 degree angle (total of 8 neighbors for each word) & iteratively learned via GNN segment Word-embeddings (RoBERTa/LayoutLMv3) + width + height n/a undirected Relative distance + direction GraphSAGE 1 n/a
94 Fully connected word / n/a directed / Token paths predicted (binary classification) 1 Greedy search
96 k-NN (k=36) segment SBERT-embeddings + normalized coordinates + Swin-Transformer-embeddings n/a undirected Relative coordinates GAT + multi-head attention 12 Sequence labeling layer
G. Grid-based approaches
Table ST6. Analysis of grid-based approaches
# Granularity Grid size Grid features Method of KIE
Height Width Feature vector size Content elements Background elements
2 character Image height Image width 54 1-hot encoded characters all-zero vector Semantic Segmentation, Bounding Box Regression (VGG-type)
3 word Image height Image width 768 BERT-embeddings all-zero vector Semantic Segmentation, Bounding Box Regression (VGG-type)
4 word 512 512 35 Character-based encoding of word + 1-hot encoded neighbor + RGB-channels all-zero vector Softmax-regression
8 word Image height Image width 4x128x103 1-hot encoded characters of N-grams discarded Custom (Attend to memory bank, Copy out attended information, parse into output)
12 character Image height Image width 257 1-hot encoded indexes of characters all-zero vector Semantic Segmentation (Coupled U-Net)
33 word Image height Image width n/a Word embeddings (Word2Vec/ Fasttext) + empty RGB-channels RGB-channels Semantic Segmentation (UNet)
40 word Image height / 8 Image width / 8 256 Word embeddings (BERT/ RoBERTa/ LayoutLM) + ResNet18-FPN-embeddings all-zero vector Binary classifier followed by set of binary classifiers, Auxiliary Semantic Segmentation
48 token vertical range of bboxes vertical range of bboxes 256 ResNet18-embeddings + horizontal and vertical relative coordinates all-zero vector BiLSTM
66 character 336 256 768 Static BERT-embeddings all-zero vector Semantic Segmentation (VGG-type), Line Item Detection
76 word Image height / 8 Image width / 8 768 PhoBERT-embeddings + Swin Transformer-embeddings all-zero vector Multi-CNN + classification layer
H. Evaluation
Table ST7. Analysis of evaluation methodologies
# Evaluation setup Granularity Unknowns performance presented Benchmarks against related work Metrics KIE datasets F1 scores
Element-based String-based Overall performance presented Field-level performance presented CORD FUNSD SROIE
1 Precision, Recall, F1 Custom
2 dist Custom
3 dist Custom
4 n/a n/a Accuracy Custom
5 meA, meP, meR, meF Ticket
6 n/a n/a F1 Custom
7 Precision, Recall, F1 Custom
8 n/a n/a Accuracy Custom
9 n/a n/a Precision, Recall, F1 Custom
10 n/a n/a F1 FUNSD, IEHHR 0,6400
11 Accuracy SROIE*, Custom
12 mean IOU, pixel accuracy, F1 Custom
13 Accuracy, F1 SROIE, Custom
14 n/a n/a F1 SROIE, Custom 0,9440
15 Precision, Recall, F1 SROIE*, Custom
16 n/a n/a ROC AUC, F1 Custom
17 n/a n/a F1 Custom
18 dist Custom
19 Hit@1, Hit@2, Hit@5 FUNSD, Custom
20 F1 Custom
21 Precision, Recall, F1 FUNSD, SROIE 0,7927 0,9524
22 mEP, mER, mEF SROIE, Ticket, Custom
23 n/a n/a F1 SROIE, Custom 0,9618
24 n/a n/a Precision, Recall, F1 CORD, FUNSD, Kleister-NDA 0,9699 0,8455
25 n/a n/a Accuracy, Precision, Recall, F1 SROIE, Custom 0,8920
26 Precision, Recall, F1 FUNSD 0,6652
27 F1 CORD, Kleister-NDA, Kleister-Charity, SROIE, SROIE* 0,9441 0,9817
28 F1 CORD, FUNSD 0,9894 0,8796
29 Precision, Recall Custom
30 n/a n/a F1 Custom
31 F1, nTED, A/B test Custom
32 F1 CORD, CORD*, FUNSD, Custom 0,9250 0,7160
33 dist, Field Accuracy Rate Custom
34 Accuracy Custom
35 Precision, Recall, F1 Custom
36 F1 FUNSD, Custom 0,5722
37 n/a n/a Precision, Recall, F1 FUNSD 0,8514
38 n/a n/a F1 FUNSD 0,8366
39 n/a n/a Precision, Recall, F1 EPHOIE, FUNSD, SROIE 0,8309 0,9688
40 F1 SROIE, Custom 0,9640
41 n/a n/a F1 CORD, Custom 0,9571
42 n/a n/a Accuracy, Precision, Recall, F1 CORD, FUNSD, SROIE 0,9477 0,8109 0,9303
43 F1 CORD, SROIE 0,9633 0,9810
44 Accuracy, F1 MARG, NIST, Tobacco, Custom
45 n/a n/a F1 EPHOIE, FUNSD, SROIE 0,8133 0,9657
46 F1 Custom
47 n/a n/a F1 EPHOIE, SROIE 0,9612
48 n/a n/a F1 EPHOIE, SROIE, Custom 0,9654
49 F1 CORD, FUNSD, Kleister-NDA, SROIE 0,9601 0,8420 0,9781
50 n/a n/a Precision, Recall, F1 FUNSD, Custom 0,6596
51 F1 Custom
52 F1 CORD, EPHOIE, FUNSD*, LastDoc4000, SROIE, XFUND* 0,9684 0,9790
53 Precision, Recall, F1 CORD, FUNSD*, SROIE, SROIE* 0,9745 0,9821
54 n/a n/a Accuracy, Recall SROIE
55 n/a n/a Precision, Recall, F1 FUNSD 0,8652
56 n/a n/a F1 FUNSD, XFUND 0,8335
57 n/a n/a Precision, Recall, F1 CORD, FUNSD, SROIE* 0,9728 0,8452
58 n/a n/a F1 CORD, FUNSD 0,9746 0,9208
59 TED-Accuracy, F1 CORD, Ticket, Custom 0,8410
60 n/a n/a F1 CORD, FUNSD, Custom 0,9728 0,8469
61 n/a n/a F1 FUNSD, CORD 0,9700 0,4610
62 n/a n/a F1 UTD, Custom
63 n/a n/a F1 CORD, FUNSD, SROIE, Kleister-NDA 0,9721 0,9312 0,9755
64 n/a n/a Precision, Recall, F1 CORD, FUNSD, EPHOIE, XFUND 0,9607 0,8415
65 n/a n/a F1 FUNSD, CORD, SROIE 0,9738 0,8649 0,9791
66 Tag Error Rate Custom
67 n/a n/a F1 Custom
68 n/a n/a Precision, Recall, F1 FUNSD, Custom 0,8346
69 n/a n/a F1 SROIE 0,9794
70 n/a n/a F1 SROIE 0,9822
71 Precision, Recall, F1 CORD, FUNSD, SROIE 0,9575 0,8345 0,9740
72 GAnTED, F1 FUNSD, NAF, IAM NER 0,6500
73 n/a n/a F1 CORD, WildReceipt 0,9541
74 Precision, Recall, F1 EPHOIE, SCID, SROIE 0,9693
75 F1, Document Accuracy Rate CORD, SROIE, Custom 0,8450 0,8730
76 F1 VAIPE
77 n/a n/a F1 CORD, WildReceipt 0,9665
78 F1 FUNSD 0,8225
79 n/a n/a F1 FUNSD, CORD, Kleister-NDA 0,9783 0,9101
80 Precision, Recall, F1 SROIE 0,9363
81 F1 CORD, FUNSD, SROIE 0,9412 0,9032 0,9788
82 F1 FUNSD, XFUND 0,8079
83 n/a n/a F1 SROIE, POIE 0,8587
84 n/a n/a Precision, Recall, F1 CORD, FUNSD, SROIE, Custom 0,9737 0,8635 0,9831
85 n/a n/a Precision, Recall, F1 CORD, FUNSD, XFUND 0,9775 0,9439
86 F1 CORD, FUNSD 0,9820 0,8400
87 n/a n/a F1 CORD, FUNSD 0,$97$ 0,9286
88 n/a n/a F1 CORD, FUNSD 0,9749 0,9376
89 F1 CORD, DeepForm, FUNSD, Kleister-Charity, PWC 0,9758 0,9162
90 n/a n/a F1 CORD, FUNSD, SROIE 0,9719 0,9320 0,9727
91 n/a n/a Precision, Recall, F1 CORD, FUNSD 0,9693 0,8877
92 n/a n/a F1 CORD, FUNSD, SIBR, XFUND 0,9565 0,9112
93 1-NED FUNSD
94 F1 CORD-r, FUNSD-r
95 F1 FUNSD, SROIE 0,7527 0,9610
96 n/a n/a F1 CORD, FUNSD, SROIE 0,9693 0,8795 0,9845